### v1

# An Extensive Study on D2C: Overfitting Remediation in Deep Learning Using a Decentralized Approach

Abstract
Overfitting remains a significant challenge in deep learning, often arising from data outliers, noise, and limited training data. To address this, we propose Divide2Conquer (D2C), a novel technique to mitigate overfitting. D2C partitions the training data into multiple subsets and trains identical models independently on each subset. To balance model generalization and subset-specific learning, the model parameters are periodically aggregated and averaged during training. This process enables the learning of robust patterns while minimizing the influence of outliers and noise. Empirical evaluations on benchmark datasets across diverse deep-learning tasks demonstrate that D2C significantly enhances generalization performance, particularly with larger datasets. Our analysis includes evaluations of decision boundaries, loss curves, and other performance metrics, highlighting D2C’s effectiveness both as a standalone technique and in combination with other overfitting reduction methods. We further provide a rigorous mathematical justification for D2C’s underlying principles and examine its applicability across multiple domains. Finally, we explore the trade-offs associated with D2C and propose strategies to address them, offering a holistic view of its strengths and limitations. This study establishes D2C as a versatile and effective approach to combating overfitting in deep learning. Our codes are publicly available at: https://github.com/Saiful185/Divide2Conquer.
Index Terms:
Classification, Deep Learning, Hyperparameter, Overfitting, Regularization, Variance.
I Introduction
Deep Learning models often do well on the data they train on, but the performance drops massively on unseen data from different distributions. This phenomenon is known as Overfitting. Many methods have been proposed to reduce overfitting over the years [10]. Early Stopping [27] is one of the most intuitive methods for reducing overfitting. However, it limits the model’s learning potential [28]. Network reduction reduces variance by simplifying the model, but it restricts learning of complex features [29]. Data augmentation is a common approach to mitigate overfitting in deep learning applications [32]. However, choosing suitable augmentation techniques for specific datasets can be challenging, and obtaining additional training data often requires substantial effort. Regularization is a widely used technique to reduce overfitting by penalizing model dependency on training data. Dropout, the most relevant regularization method in deep learning, reduces overfitting by randomly dropping connections between neural network layers [4]. However, even dropout can sometimes degrade the performance of a model due to internal covariance shifts [18]. Eventually, despite employing these techniques, the model still has access to the entire training data, and the robust neural network structures find their ways to fit on the training set a bit too much. This is where our motivation comes from. It can be a good idea to not let the neural network train on the entire dataset in the first place. Intuitively, we should combine multiple models that train on different portions of the data so that the model can not familiarize itself with all of it at once, and a consensus can be achieved by focusing on the general patterns observed in different portions of the data.
We propose a novel method, Divide2Conquer (D2C), which proposes dividing the training data into multiple subsets and training a model(all having the same architecture & hyperparameters) on each subset. A weighted averaging of all the parameter weights from the instance model is performed after a certain number of epochs, and the averaged weights are shared back to each of the models (we denote them as edge models in this study). Then, the whole process is repeated for several global epochs.
D2C is loosely inspired by Federated Optimization[2], which involves training identical models on local devices (using the data available on those devices) and periodically aggregating and averaging their trained parameters on a central server. This process ensures safe access to data and preserves privacy, as it only shares the weights, not the data itself. However, unlike D2C, federated learning is not inherently designed to enhance model generalization.
The D2C method also has similarities with Bagging[23] in the form of using subsets of data. However, in bagging (Bootstrap Aggregating), the data is not completely separated for different models as we proposed. Instead, the technique involves creating multiple subsets of the original dataset by randomly sampling with replacement. This means that some datapoints may be repeated in a subset while others may be excluded altogether. Also, in bagging, the aggregation is done by combining the output probabilities, not the model weights.
Our experiments encompass multiple datasets across different domains, and each case suggests that employing the D2C method can reduce overfitting significantly. The primary comparison to evaluate the method was between the performances of the base Neural Network architecture used for the Edge Models using the entire training data and the performance of our models using the Divide2Conquer method. We summarize our contributions through this study below:
  • We introduce a new method, D2C, that helps in reducing overfitting significantly while being conceptually simple and easy to implement.
  • We present a mathematical justification for the hypothesis that D2C mitigates overfitting by rigorously demonstrating its capacity to reduce model variance. This formal analysis underscores D2C’s potential to improve model robustness across diverse datasets.
  • D2C can be applied on top of other data augmentation and regularization techniques, and our experiments across various datasets and tasks show that this results in a clear improvement in the model’s generalization ability.
  • We extensively tune the hyperparameters introduced by this method and report the findings, providing important directions for future applications.
  • We also delve into the trade-offs associated with D2C, including its computational overhead and potential diminishing returns for smaller datasets. We explore strategies to mitigate these challenges and provide a comprehensive understanding of D2C’s strengths and weaknesses.
The paper is organized as follows. In section II, we discuss relevant literature. We establish the theoretical justification behind our hypothesis in section III. In sections IV and V, we discuss our methodology and lay out the experimental specifications. In section VI, we evaluate our approach through empirical experiments on multiple datasets and analyze the results. We discuss the limitations of D2C in section VII. Section VIII contains concluding remarks.
II Related Works
Several authors have addressed the issue of overfitting while performing various tasks. M. Cogswell et al. proposed a regularizer called DeCov, which helps reduce overfitting in deep neural networks by Decorrelating Representations[5].
Dropout [4] was proposed by Srivastava et al. back in 2014, and since then, this technique has been extensively used in very complex neural network architectures successfully. It was indeed an outstanding contribution, specifically for deep learning-based models. Batch Normalization [8] proposed by Ioffe and Sergei primarily focuses on better convergence and somewhat contributes to reducing overfitting. J. Kolluri and V.K Kotte came up with L¼ regularization to solve the problems faced by L1 and L2 regularization techniques [15].
Ensembles are also often used to improve generalization. The ensemble-based ELVD model [3] managed to outperform the traditional VGGNet and DropoutNet models in terms of reducing overfitting. Zehong Zeng et al. came up with an ensemble framework that incorporates several techniques to prevent overfitting [9]. Min-Gu Kim et al. also proposed parallel ensemble networks to reduce overfitting in ECG data and prevent the degradation of generalization performance as the training progresses [1].
We implemented a method based on federated optimization preliminarily for facial expression recognition using FedNet[17]. This model achieves excellent results in terms of generalization on both CK+ and FER-2013 datasets. The K-Fold-Cross-Validation method for evaluation proposed by Korjus K. et al. [13] has some similarities to our proposed method in the sense that it also creates subsets by partitioning the dataset and makes use of the whole training data. However, this method is different from ours since the erroneous samples still make it to the training phase (K-1/K)-th of the time in K-Fold CV. The proposed D2C method partitions data and the whole training process is partitioned too, which means none of the subsets is fed with a particular outlier while training bar one, minimizing the impact of these erroneous samples.
Overfitting remains a prevalent issue in supervised machine learning despite methods like Early Stopping, Network Reduction, Training Set Expansion, Regularization, and Dropout being effective [10]. Through D2C, we build an approach that focuses on achieving even better generalization and can be implemented on top of other overfitting-reducing techniques.
III Theoretical Framework
III-A The Hypothesis
Overfitting happens when a model fits too well on the data that it trains on. It effectively captures and memorizes even the random characteristics from the training data, randomness that would be insignificant in real-world applications. The presence of outliers and noisy data is a fundamental reason behind overfitting. In this study, we formulate a method to minimize the effect of these outliers and noisy data.
The contribution of outliers/noise can be minimized by dividing the training data into multiple shards and training each shard separately. That way, each data point will only occur once in one of the several data shards. The representative samples would be close to each other in the feature space and would be present in each of the data shards more or less uniformly. However, the individual outliers/noise would only be able to impact one of the training processes. Aggregation and averaging can be done periodically to combine all the models. In the best-case scenario, due to averaging, the effect of the individual outliers/noise would be reduced by a factor of N, where N is the number of data shards. The key factors behind our hypothesis and considerations regarding this method being able to address overfitting are discussed as follows:
1. Weighted averaging of parameters helps in combining the knowledge learned from different subsets. However, extreme parameter updates driven by noise or outliers in individual models are moderated when averaging across multiple models. If a model encounters a noisy sample that deviates significantly from the overall data distribution, it might adjust its parameters excessively to fit that sample, leading to overfitting. However, if the training data is divided into subsets and trained separately, even if one model’s parameters are influenced by noisy or outlier samples, their impact is diluted when combined with the parameter weights of other models during averaging. So, the averaged parameters should reflect a more balanced representation of the underlying data distribution, mitigating the influence of individual noisy or outlier samples.
2. Averaging the parameter weights is particularly helpful in combating overfitting since it has a regularizing effect. Averaging weights regularizes the model at the parameter level. It moderates extreme weight updates driven by noise or outliers in individual models, leading to more stable and generalizable parameter values. This helps prevent overfitting by discouraging models from fitting to the noise or idiosyncrasies present in their training subsets. By averaging parameter weights, the central model converges to a shared solution that reflects the collective knowledge learned from different perspectives. This consensus learning approach encourages models to learn generalizable patterns and reduces reliance on individual model predictions that are prone to overfitting.
3. Averaging parameters smooths out the decision boundary learned by individual models. Extreme or noisy parameter updates that result in sharp or jagged decision boundaries in individual models are likely to be moderated and smoothed when combined through averaging, consequently helping in generalizing well to unseen data and reducing overfitting.
4. The optimal hyperparameters, such as the number of subsets and epochs before each round of central averaging vary by dataset and domain. The appropriate subset count depends on dataset size and class distribution. Suppose the edge model is complex or the subset count is too high. In that case, the overly reduced size and number of samples per class in each subset can result in significant overfitting in edge models, which in turn can degrade overall performance even after central averaging. This is similar to what happens in federated learning when the client data is minimal, as pointed out by Zhang et al. [34]. When sufficient data exists per class, more subsets can be beneficial, as noise or outliers are diluted by a factor of up to N (number of training subsets).
In summary, D2C helps to reduce overfitting by mitigating the effects of noise and outliers. However, dividing the training set into too many subsets can lead to insufficient data per subset, which can hamper the overall performance.
III-B Mathematical Intuition
To lay out the underlying intuition behind our hypothesis, we must first define overfitting formally. We follow the groundwork laid out by Ghojogh et al. [23]. Let us consider a Dataset $D$ with training set $T$ and test set $R$ so that,
$T\cup R=D$ (1)
$T\cap R=\emptyset$ (2)
Let, $T$ consist of datapoints $\{(x_{i},y_{i})\}_{i=1}^{n}$ and test set $R$ consist the datapoints $\{(x_{i},y_{i})\}_{i=1}^{m}$. Also, let $f$ be the function that maps each datapoint in the training set to its corresponding output label(true observations), i.e., $f(x_{i})=f_{i}$, where $f_{i}$ is the true corresponding output label for $x_{i}$. This model is unknown and we wish to train a model that replicates $f$ as closely as possible. However, both the ground truth model and the true labels are unknown to us. Realistically, there is always some noise in the training dataset. We assume that the training output is corrupted with some additive noise $\epsilon_{i}$. So, we can say,
$y_{i}=f_{i}+\epsilon_{i}$ (3)
where $\epsilon_{i}$ is some Gaussian noise, $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. Since the mean of the distribution is zero we can say that $\mathop{{}\mathbb{E}}(\epsilon_{i})=0$, and $Var(\epsilon_{i})=\sigma^{2}$. As the variance of a random variable $x$ can be denoted as $Var(X)=\mathop{{}\mathbb{E}}(x^{2})-[\mathop{{}\mathbb{E}}(x)]^{2}$ (Appendix A), we have,
$\mathop{{}\mathbb{E}}(\epsilon_{i}^{2})=Var(\epsilon_{i})+[\mathop{{}\mathbb{E}}(\epsilon_{i})]^{2}\Rightarrow\mathop{{}\mathbb{E}}(\epsilon_{i}^{2})=\sigma^{2}+0=\sigma^{2}$ (4)
Our assumption is that the input of the training dataset ${\{{x}_{i}}\}_{i=1}^{n}$ and their noise-added outputs ${\{{y}_{i}}\}_{i=1}^{n}$ are available to us, and our goal is to estimate the true model $f$ by a model $\hat{f}$ in order to predict the labels ${\{{y}_{i}}\}_{i=1}^{n}$ for the input data ${\{{x}_{i}}\}_{i=1}^{n}$. Let the estimated observations be ${\{\hat{y}_{i}}\}_{i=1}^{n}$. It is preferable that ${\{\hat{y}_{i}}\}_{i=1}^{n}$ are as close as possible to ${\{{y}_{i}}\}_{i=1}^{n}$. Mathematically, we can consider this as a minimization problem looking to minimize some error function such as MSE (Mean squared error). The MSE of estimated outputs from the trained model with respect to the dataset outputs are: $\textbf{MSE}=\frac{1}{n}\sum_{i=1}^{n}(\hat{y}_{i}-y_{i})^{2}=\mathop{{}\mathbb{E}}((\hat{y}_{i}-y_{i})^{2})$. Let us consider a single arbitrary datapoint $(x_{0},y_{0})$ and the corresponding prediction of the trained model $\hat{f}$ for the input $x_{0}$ is $\hat{y}_{0}$. Ghojogh et al. [23] showed that if the sample instance $(x_{0},y_{0})\notin T$ i.e., it belongs to the test set $R$, MSE of the predicted label can be expressed as -
$\textbf{MSE}=\mathop{{}\mathbb{E}}((\hat{y}_{0}-y_{0})^{2})=\mathop{{}\mathbb{E}}((\hat{y}_{0}-f_{0})^{2})+\sigma^{2}$ (5)
Here we can observe from the first term of the R.H.S. of the equation that the MSE of the estimation of output labels with respect to the dataset labels can actually be represented by the MSE of the estimation with respect to the true uncorrupted labels ${\{{f}_{i}}\}_{i=1}^{n}$ plus some effect of the noise that exists in the dataset. Taking Monte-Carlo approximation [24] of the expectation terms in Eq. (5), we have:
$\frac{1}{m}\sum_{i=1}^{m}{(\hat{y}_{i}-y_{i})^{2}}=\frac{1}{m}\sum_{i=1}^{m}{(\hat{y}_{i}-f_{i})^{2}}+\sigma^{2}$
$\implies\sum_{i=1}^{m}{(\hat{y}_{i}-y_{i})^{2}}=\sum_{i=1}^{m}{(\hat{y}_{i}-f_{i})^{2}}+m\sigma^{2}$ (6)
Here the term $\sum_{i=1}^{m}{(\hat{y}_{i}-y_{i})^{2}}$ is the total error between the model predictions and the dataset labels which we can call the empirical observed error, $e$. On the other hand, $\sum_{i=1}^{m}{(\hat{y}_{i}-f_{i})^{2}}$ represents the total error between prediction labels and the true unknown labels, namely $E$. Hence, $e=E+m\sigma^{2}$. Thus, the observed error truly reflects the true error because the term $m\sigma^{2}$ remains constant, which is the theoretical justification for evaluating model performance using unseen test data. But, as Ghojogh et al. [23] showed, if $(x_{0},y_{0})\in T$, deriving the M.S.E yields:
$\textbf{MSE}=\mathop{{}\mathbb{E}}((\hat{y}_{0}-y_{0})^{2})=\mathop{{}\mathbb{E}}((\hat{y}_{0}-f_{0})^{2})+\sigma^{2}-2\sigma^{2}\mathop{{}\mathbb{E}}(\frac{\partial\hat{y}_{0}}{\partial y_{0}})$ (7)
Using the same approximation as Eq.(6) on the training set $T$, we get:
$\frac{1}{n}\sum_{i=1}^{n}{(\hat{y}_{i}-y_{i})^{2}}=\frac{1}{n}\sum_{i=1}^{n}{(\hat{y}_{i}-f_{i})^{2}}+\sigma^{2}-2\sigma^{2}\frac{1}{n}\sum_{i=1}^{n}\frac{\partial\hat{y}_{i}}{\partial y_{i}}$
$\implies\sum_{i=1}^{n}{(\hat{y}_{i}-y_{i})^{2}}=\sum_{i=1}^{n}{(\hat{y}_{i}-f_{i})^{2}}+n\sigma^{2}-2\sigma^{2}\sum_{i=1}^{n}\frac{\partial\hat{y}_{i}}{\partial y_{i}}$ (8)
From Eq.(8) we now understand while the model is training the observed empirical error calculated on the training dataset i.e, training error is not a clear representation of the true error because now $e=E+n\sigma^{2}-2\sigma^{2}\sum_{i=1}^{n}\frac{\partial\hat{y}_{0}}{\partial y_{0}}$. Hence, even if we can obtain a small value of $e$, the term $2\sigma^{2}\sum_{i=1}^{n}\frac{\partial\hat{y}_{0}}{\partial y_{0}}$ can grow large as the training progresses and hide the true value of a substantial and large true error $E$ (Figure 1). As we can see in the figure, the model starts to overfit as empirical error becomes no longer a true representation of true error when the corresponding complexity term grows and dominates the expression in Eq.(8) for empirical error. We can conclude from this analysis that the final term in Eq.(7) is a measure of the overfitting of the model. Which is also known as the complexity of a model. Upon a closer look at this final term, we can derive some more interesting insight. $\frac{\partial\hat{y}_{0}}{\partial y_{0}}$ is essentially the rate of change of the predicted label $\hat{y}_{i}$ for the input $x_{i}$ with respect to the change in dataset label $y_{i}$. It makes sense that prediction labels would vary if the corresponding output in the dataset varies as the sample is considered from within the training set. It also reflects how dependent the model is on the training set, as the term grows when the predicted output varies too much with respect to changes in the training samples. This fits the classical definition of overfitting i.e., fitting too tightly to the training data and thus performing way worse when evaluated with actual test data. Evaluation on test data accurately represents the true error from Eq.(6). Figure 2 depicts a visualization of this idea.
Figure 1: When the corresponding complexity term grows and dominates the expression in Eq.(8) for empirical error, the model starts to overfit as empirical error becomes no longer a true representation of true error.
Figure 2: A visualization of (a) Underfitting, (b) A good fit, and (c) Overfitting scenarios, illustrating how the model changes if an arbitrary training data point $(x,y)$ is shifted to $(p,q)$ by plotting the trained model before (solid line) and after (dashed line) the shift.
Statistical analysis shows that the distinction between an underfitting model, an overfitting one, and a good fit can be reflected by their bias and variance with respect to the training labels. Ideally, a model with low variance and low bias is the most desirable and usually deemed a good fit. However, during the training phase, sometimes we see a model has low variance and high bias which indicates the model predictions are not good as well as being all over the place, this is the indication of an underfitting model. It usually happens when the model is too simple to capture the complexity of the data or has not been trained for long enough to reach convergence. On the other hand, we find a model which overfits has high variance and low bias. This is because the model becomes overly complex accommodating for all the outliers in the training data and fails to generalize on unseen test data. Figure 2(a) and 2(c) show how an underfitting and overfitting model shows low and high variance respectively. Figure 2 is a visualization of (a) Underfitting, (b) A good fit, and (c) Overfitting scenarios. The figures illustrate how the model changes if an arbitrary training data point $(x,y)$ is shifted to $(p,q)$ by plotting the trained model before and after the shift. In (a) the underfitted model the change is very minimal. In the (b) Good fit model the change is also relatively little, but in the (c) Overfitted model there is a large shift in the neighboring region of that data point, which reflects the over-dependence on training data.
To see the impact of D2C we consider the training set $T$ divided into $k$ equal partitions, namely $\lvert T_{1}\rvert=\lvert T_{2}\rvert=...=\lvert T_{k}\rvert=\lvert T\rvert/k$. Then, we train the model $\hat{f}_{j}$ using the $j$-th data chunk, where $j\in\{1,..,k\}$. As a result, we have $k$ separately trained models. Finally, we aggregate these edge models into one central model $\hat{f}$. This process is repeated for $p$ global epochs. For the sake of simplicity, we consider each $\hat{f}_{j}$ is a regression model with weight matrix $\boldsymbol{W_{j}}=\{w_{j1},...,w_{jn}\}$ and bias $b_{j}$. For an input vector $\boldsymbol{x}=\{x_{1},...,x_{n}\}$ the prediction of the edge model on the $t$-th epoch can be described as: $\hat{f}_{j}(\boldsymbol{x})=\boldsymbol{W_{j}}^{t}\boldsymbol{x}^{T}+b_{j}$. In the case of a single global epoch, we now wish to aggregate the weight matrices creating the central weight matrix $\boldsymbol{\bar{W}}=\{\bar{w}_{1},...,\bar{w}_{n}\}$. Here, the weight matrix for the $(t+1)$-th epoch is:
$\boldsymbol{\bar{W}}^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\boldsymbol{W_{j}}^{t}\textrm{ and }b^{t+1}=\frac{1}{k}\sum_{j=1}^{k}{b^{t}_{j}}$ (9)
Hence, the output of the central model at $(t+1)$-th epoch:
$\hat{f}(\boldsymbol{x})^{t+1}=\boldsymbol{\bar{W}}^{t+1}\boldsymbol{x}^{T}+b^{t+1}$
$\implies\hat{f}(\boldsymbol{x})^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\boldsymbol{W_{j}}^{t}\boldsymbol{x}^{T}+\frac{1}{k}\sum_{j=1}^{k}{b^{t}_{j}}$
$\implies\hat{f}(\boldsymbol{x})^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\left(\boldsymbol{W_{j}}^{t}\boldsymbol{x}^{T}+{b^{t}_{j}}\right)$
$\implies\hat{f}(\boldsymbol{x})^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\hat{f}(\boldsymbol{x})^{t}_{j}$ (10)
Let $e_{jt}$ denote the empirical error of the $j$-th model while estimating some arbitrary instance on the $t$-th global epoch. We assume this error to be a random variable with normal distribution having mean zero, i.e., $e_{jt}\sim\mathcal{N}(0,s)$. Hence, $\mathop{{}\mathbb{E}}(e_{jt})=0$ and $Var(e_{jt})=s$. Let us denote the estimation of the product of two trained models using two different data chunks by $c$, i.e, $\mathop{{}\mathbb{E}}(e_{jt}\times e_{lt})=c$. So, we have:
$\mathop{{}\mathbb{E}}(e_{jt}^{2})=Var(e_{jt})+[\mathop{{}\mathbb{E}}(e_{jt})]^{2}\Rightarrow\mathop{{}\mathbb{E}}(e_{jt}^{2})=s+0=s$ (11)
Again for $j,l\in\{1,...,k\}$, where $j\neq l$, we have:
$Cov(e_{jt},e_{lt})=\mathop{{}\mathbb{E}}(e_{jt}\times e_{lt})-\mathop{{}\mathbb{E}}(e_{jt})\mathop{{}\mathbb{E}}(e_{lt})$
$\implies Cov(e_{jt},e_{lt})=c-0\times 0=c$
$\implies Cov(\hat{f}(\boldsymbol{x})^{t}_{j},\hat{f}(\boldsymbol{x})^{t}_{l})=c-0\times 0=c$ (12)
From Eq.(10) and Eq.(12), we can deduce:
$Var(\hat{f}(\boldsymbol{x})^{t+1})$ $=\frac{1}{k^{2}}Var\left(\sum_{j=1}^{k}\hat{f}(\boldsymbol{x})^{t}_{j}\right)$ (13)
$=\frac{1}{k^{2}}\sum_{j=1}^{k}Var\left(\hat{f}(\boldsymbol{x})^{t}_{j}\right)+\frac{1}{k^{2}}\sum_{i=1,l=1}^{k}$
$\sum_{i\neq l}^{k}Cov\left(\hat{f}(\boldsymbol{x})^{t}_{i},\hat{f}(\boldsymbol{x})^{t}_{l}\right)$
$=\frac{1}{k^{2}}sk+\frac{1}{k^{2}}ck(k-1)=\frac{s+c(k-1)}{k}$
Relevant proofs related to the derivation of Eq. 12 and 13 are provided in Appendix A. Eq.(13) provides us with an interesting insight. If the two edge models are highly correlated i.e., their predictions and empirical error are very close to each other we can assume $s\approx c$. Hence, the variance of the central model at $(t+1)$-th epoch can be reduced to: $Var(\hat{f}(\boldsymbol{x})^{t+1})\approx(s+s(k-1))/k\approx s$. On the other hand, if the two models are very dissimilar the same expression gives us: $Var(\hat{f}(\boldsymbol{x})^{t+1})\approx(s+0(k-1))/k\approx s/k$. As $Var(e_{jt})=Var(\hat{f}(\boldsymbol{x})^{t})=s$, each of the edge models at epoch $t$ has a variance of prediction $s$, which means if the edge models at epoch $t$ are completely different, the variation of estimate decreases by a factor of $k$ on epoch $(t+1)$. Consequently, if the edge models are identically trained the variance remains almost unchanged. This result is similar to the one proposed by Breiman et al. [25] and Bühlmann et al. [26] through the idea of ’Bagging’, which is a meta-algorithm aggregating multiple model outputs. The analysis clearly shows that our approach does not at the least worsen the problem of overfitting in this scenario but rather improves the central model on each epoch if certain conditions are met. However, in reality, the edge models are neither completely different nor exactly the same due to the relative distribution of the data subsets and their shared model architecture. So, the degree to which D2C ameliorates overfitting depends on the model architecture and the correlation among the edge models. However, it is easy to see that if the training data is contaminated by outliers and random noise, the learned parameters from the different portions of the data are more likely to be different from each other. This is because the individual outliers and noise are not highly correlated, hence distributing them among the subsets likely results in significantly different samples in the subsets, making the subsequent models different from each other. Our results show that repeating the process in every global epoch reduces overfitting in most cases. A limitation of this analysis is that deep neural networks are mostly black boxes and their architectures vary tremendously, which makes establishing a generalized mathematical representation of them extremely difficult. Hence, our assumption of a simple regression model is not an exact reflection of the actual implementation. Despite that, it describes an ideal and simple scenario where the proposed method should improve the model. Moreover, it also gave us the insight that this approach can be used in conjunction with other overfitting reduction techniques such as Dropouts(Srivastava et al., 2014) [4]. If dropouts are stacked on top of our method in every iteration of local training, the neurons are randomly removed with some probability $p$(usually $p=0.5$). This introduction of randomness ensures that edge models are different from each other increasing the chances of improvement when the models are aggregated.
IV Proposed Methodology
IV-A Creating Training Subsets
To train multiple Neural Networks parallelly using different subsets of the data, we need to divide the whole training set into multiple shards, and we also need multiple identical models that will be used to train the different subsets. These models will be denoted as Edge models in this paper. At first, the training set is randomly shuffled and then divided into multiple subsets, maintaining consistent class distribution. Each subset is then fed into a separate edge model with identical neural network architecture and trainable parameters to enable parameter averaging.
IV-B Training & Averaging of Trained Parameters
Instead of one training loop, the implementation of our method requires two loops: an inner loop for subset training which is nested within an outer loop that performs central averaging. A central model with the same architecture as the edge models is first initialized. In total, we have one Central Model and $N$ Edge models, all of which share an identical architecture, where $N$ is the number of subsets we have divided the training set into. Initially, these identical model objects are built and loaded for each training subset. Within the inner loop, each edge model is iteratively trained on its subset for a specified number of local epochs. Then, the edge models’ parameters are scaled based on their data fraction and added to a local weight list. In the outer loop, these scaled parameters are averaged, updating the central model’s parameters. Each edge model’s weights are then reinitialized to match the central model, completing a global training epoch. This process repeats over several global epochs, with weighted parameter averaging to account for the data fraction of each subset. The whole process is summarized in Figure 3, and the pseudo-code to implement D2C is provided in Algorithm 1.
Algorithm 1 Divide2Conquer
  Input: Training dataset $D$, Number of subsets $N$, Local epochs $E$, Global epochs $E_{global}$, Batch size $B$, Learning rate $lr$
  Output: Central model $M_{c}$ with aggregated weights
  function Divide2Conquer($D$, $N$, $E$, $E_{global}$, $B$, $lr$)
        Shuffle the dataset $D$
        Divide $D$ into $N$ subsets: $\{D_{1},D_{2},\dots,D_{N}\}$, maintaining class distributions
        Initialize central model $M_{c}$ with parameter vector $\theta_{c}$
  for $i=1$ to $N$ do
           Initialize edge model $M_{i}$ with $\theta_{i}=\theta_{c}$
           Set scaling factor $s_{i}=\frac{|D_{i}|}{|D|}$
  end for
  for $g=1$ to $E_{global}$ do
           Initialize weight accumulator $W_{central}=0$
     for $i=1$ to $N$ do
                  Set $\theta_{i}=\theta_{c}$
        for $e=1$ to $E$ do
                         Train $M_{i}$ on $D_{i}$ using batch size $B$ and learning rate $lr$
        end for
                  Obtain updated weights $\theta_{i}$ from $M_{i}$
                  Scale weights: $\theta_{i}=s_{i}\cdot\theta_{i}$
                  Accumulate scaled weights into $W_{central}$: $W_{central}=W_{central}+\theta_{i}$
     end for
           Compute averaged central weights:
$\theta_{c}=\frac{1}{\sum_{i=1}^{N}s_{i}}W_{central}$
           Update central model $M_{c}$ with new weights $\theta_{c}$
  end for
  return central model $M_{c}$
  end function
IV-C Global Hyperparameter Tuning
Tuning the overall model consists of two stages. Firstly, the edge model and its hyperparameters should be tuned using a subset/entire training set and the validation set to find a suitable model to train each training subset. Next, that base model is utilized to tune the Central model. The new global hyperparameters that need tuning, in this case, are the Number of Subsets of the Training Set and the Number of Epochs before each round of Global Averaging. Testing for different values for these hyperparameters, the implementation, and the evaluation of our method are done using the best combination. The appropriate number of training subsets can be determined first by varying the number of training subsets and observing the performance metrics like test/validation accuracy, F1 score, log loss, and also the validation loss curve. The appropriate number of training epochs in each subset before the averaging is done each time is likely to be a small number $(1\sim 3)$ since the edge model can quickly overfit the training subset with a reduced size. So, it makes more sense to set a small number of epochs initially and tune the number of training subsets. Once the number of subsets is determined, the same process can be repeated by varying the number of training epochs and determining the appropriate value of it.
Figure 3: D2C at a Glance – Training on each subset, averaging the trained parameters, setting the Central model parameters to these values, and sharing the averaged parameters back to the Edge models to keep the loop running.
V Experimental Setup
V-A Datasets
For image classification tasks, we employed three widely used benchmark datasets: CIFAR-10 [14], MNIST [7], and Fashion MNIST [12]. CIFAR-10 provides a challenging image classification task with 60,000 32x32 RGB images across 10 categories. MNIST offers a simpler task with 70,000 28x28 grayscale images of handwritten digits. Fashion MNIST presents a more complex variation with 70,000 28x28 grayscale images of fashion items. To address the overfitting challenge in facial expression recognition, we utilized the FER-2013 [16] dataset, which comprises over 35,000 grayscale images of seven basic emotions. The diverse and noisy nature of this internet-sourced dataset makes it particularly suitable for evaluating generalization performance.
For audio-based tasks, we created a comprehensive dataset by combining TESS [19], CREMA-D [20], and RAVDESS [21]. This dataset encompasses 14 emotion classes, offering a diverse range of emotional expressions and recording conditions. By merging these datasets, we aim to improve model generalization by mitigating biases and limitations inherent in individual datasets.
To test our method on text data, we employed the AG News [22] dataset, which consists of approximately 127,600 news articles categorized into four balanced classes: World, Sports, Business, and Science/Technology. This dataset provides a large-scale, real-world benchmark for evaluating text classification models, particularly in the context of Big Data.
V-B Experimental Details
1. Image Classification: At first, all the images are reshaped and normalized. We applied one-hot encoding to the labels to use cross-entropy loss later. We split the training data into training and validation sets using Scikit-Learn. We then divided the training set into multiple subsets. Before constructing the model, processing each training subset into tensors, and batching them was our last step. Each of our edge models was comprised of 4 Convolutional (32 to 256 kernels progressively) and MaxPooling layer (2X2) blocks followed by one fully connected hidden layer (128 neurons). Each block contained two identical convolutional layers and one pooling layer.
2. Audio Classification: At first, we combined the paths of the files of our three datasets into a single CSV file. Then we read the WAV files as time series and applied white noise. Then we converted the 1D time series into 2D Log Mel Spectrograms to capture the audio signals’ time and frequency domain features. We treated the resulting data as 2D images, and then the remaining steps were the same as we mentioned in the case of the Image classification tasks.
3. Text Classification: In this case, we created a corpus using all the words occurring in our training dataset. Then we made a word index and word embedding for all the words in the corpus. After that, we created word sequences using that word embedding for each data instance. We applied padding to make each training data point equal in length. In the case of these 1D text sequences, the edge model we used consisted of 3 bidirectional LSTM layers(128, 64, and 32 units) and two dense layers(128 and 64 neurons) followed by the output layer. It took the embeddings as input, and we used the 100-dimensional version of GloVe from Stanford for these embeddings.
In all cases, we used a 90-10 Training–Validation split, Adam optimizer, and applied dropouts and batch normalization layers. This is important to test for D2C’s effectiveness when used on top of other methods that address overfitting. We used the predefined test sets for testing.
VI Results and Discussions
We will use loss and accuracy curves, Accuracy, F1 Score, AUC-ROC, and log loss to analyze the results and evaluate our method. We also varied the two new global hyperparameters we mentioned in the previous section: the Number of Subsets of the Training Set and the Number of Epochs before each round of Central Averaging. To refer to them concisely, we will use the variables $\boldsymbol{N}$ and $\boldsymbol{E}$ respectively in this section.
(a) No Subsets (Traditional)
(b) N = 3 (D2C)
Figure 4: Decision boundary Visualization with the Traditional approach vs. the Divide2Conquer method.
Figure 5: Decision Boundary from the edge models trained on each subset after dividing the training data into 3 subsets applying the D2C method.
VI-A Visualizing Decision Boundary
At first, we will visualize the change in decision boundaries due to applying the D2C method and analyze the findings. We used a synthetic binary classification dataset to simulate a simple decision boundary. This dataset was created using the Scikit-Learn library. Then the 240-sample large dataset was divided into a training and a test set using a 75-25 split. The neural network architecture we used for these experiments comprised three hidden layers having 100 neurons each. No regularization or data augmentation was used for decision boundary visualization, to observe the effect of D2C explicitly.
At first, we adopted the centralized, traditional approach and fed the training dataset to the sole model. After training, we got the decision boundary shown in Figure 4a.
As we can see, the decision boundary for the traditional approach is highly complex and non-linear here. The model creates intricate regions to distinguish between the two classes. The boundary closely follows individual points, resulting in a jagged, convoluted shape. The complexity of the decision boundary suggests that the model is overfitting to the training data. It is trying too hard to separate every point, including noise, rather than focusing on general patterns in the data. Evidence of overfitting is seen in how the model creates small pockets of blue in the red region and vice versa. This indicates the model has likely memorized the data rather than learning generalizable patterns. This means that this approach is likely to produce excellent results on training data. Still, it is likely to struggle with new, unseen data due to its sensitivity to specific details in the training set. The highly irregular boundary reflects poor generalization ability.
Then, to see the effect of D2C on decision boundaries, we divided our training set into multiple subsets and trained them parallelly using identical Edge Models. Figure 5 shows the decision boundaries we got from each edge model.
As we can see, the decision boundaries from the edge models are still complex, containing intricate regions following individual noisy data points. However, each of these noisy samples occurs only once in one of the subsets, affecting the corresponding edge model. So, irregularity around an individual noisy sample does not affect all the edge models but rather only one of them. Hence, after averaging the weights, the impact of the individual outliers should be diluted.
To verify, the trained weights from each subset were averaged to get the final weights for the Central Model. The resulting decision boundary is shown in Figure 4b.
We can see that the decision boundary from the central model of D2C is much simpler, dividing the feature space almost diagonally. This suggests a less flexible, more generalized model due to averaging the weights from multiple subsets. This boundary is smooth and does not follow the exact positions of individual data points as closely. The model is more generalized, focusing on the overall trend in the data rather than trying to accommodate every single datapoint, exhibiting much less overfitting compared to what we saw in the traditional approach using the same model architecture. By averaging the weights from different subsets, the model smooths out the noise in the data and prevents overfitting to specific training samples. An individual noisy sample does not affect all the edge models in D2C, and its impact is diluted through averaging. This leads the central model to create a decision boundary based on the more common patterns present in all the subsets, ensuring better generalization. Further proof of this generalization is seen in the training and test accuracies for the traditional approach and D2C using different numbers of subsets. We summarize these accuracies in Table I.
TABLE I: Performance of the traditional approach vs different values of $N$ Using the D2C method.
Model Settings Training Accuracy Test Accuracy
No Subsets 0.9944 0.6833
N=2 (2 Subsets) 0.7611 0.75
N=3 (3 Subsets) 0.7944 0.8333
N=4 (4 Subsets) 0.7889 0.8333
As we can see, we got almost perfect accuracy on the training set with the traditional approach. However, the significant drop in test accuracy (0.6833) proves that the model is overfitting to the training data. When we use D2C with 2, 3, or 4 subsets, we observe a clear improvement and a much better balance between the training and test accuracies. For all these cases, the training and test accuracies are similar, and the test accuracy for 3 or 4 subsets (0.8333) surpasses that of the traditional approach by a significant margin. This indicates that splitting the data into 3 or 4 subsets provides a good balance between learning patterns from training data and reducing overfitting to improve the model’s ability to generalize to unseen data.
Our observations prove that models using the D2C approach perform better on unseen data, as they avoid overfitting by keeping the decision boundary simple and focusing on capturing the broader, more consistent patterns in the data. So, D2C can be seen as yet another technique to reduce overfitting. However, several other techniques address overfitting. So, whether using D2C on top of other regularization and augmentation techniques (such as dropouts) improves the performance of a model is the key question in application level. In the following subsections, we will investigate this using the architectures (incorporating regularization methods like dropouts in the edge models) we described in the previous section on several datasets from different domains.
VI-B Analyzing Loss & Accuracy Curves
We will analyze the results obtained from our experiments from the 3 different domains. The loss curve is a definitive indicator of overfitting. As we saw in Eq. (6), The validation/test loss truly represents the true error of the model. So we will monitor the validation loss curves from the different experiments we performed to compare and analyze if D2C improves the generalizing performance, in the form of minimizing the validation loss as the training progresses.
Image Classification: Let us first take a look at the loss curves generated using the CIFAR-10, Fashion MNIST, MNIST, and FER2013 datasets. For now, we will keep the number of epochs before each round of central averaging, $\boldsymbol{E}$, equal to $1$. It will ensure a proper comparison between the traditional method and our proposed method while varying the number of subsets, $\boldsymbol{N}$. In our experiments, how much we varied $N$ depended on the performance (Loss curve, Accuracy) trend as we increased $N$, the number of subsets. In the case of all the datasets, we stopped varying $N$ whenever we observed a definitive performance drop with increasing N. For example, in the case of CIFAR-10, we varied $N$ in the range of $1$ to $9$.
Figure 6: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on CIFAR-10.
Figure 7: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on Fashion MNIST.
Figure 8: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on MNIST.
Figure 9: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on FER2013.
As we can see from Figure 6 and Figure 9, the validation loss keeps decreasing with each epoch for longer when using our proposed D2C method on these datasets but starts to increase after very few epochs again in the case of traditional NN implementations using the entire training set. In Figure 7, we can see that D2C with 9 subsets almost entirely eliminates the trend of an increasing loss after a few epochs. In machine learning, overfitting is marked by a model’s increasing validation loss after a certain number of training epochs, indicating that the model has started to capture noise and specific variations in the training set that do not generalize to new data. A model that reaches this overfitting point later, meaning it maintains a lower validation loss over a longer period, suggests it is better regularized or has a complexity that is well-suited to the underlying patterns in the data. A model with a delayed overfitting point continues to find generalizable patterns within the data through many epochs, highlighting that it remains in a generalization phase for longer, rather than immediately entering an overfitting phase. We can also see that, in all these cases, dividing the training data into more and more subsets delays and slows down the abrupt increase in validation loss as the training progresses. For example, in the case of CIFAR-10, the loss curve generated using the traditional approach without dividing the training data (no subsets) reaches its minimum, $0.536$, at around $16$ epochs. Then it keeps increasing and reaches $0.77$ by the end of 100 epochs. As we set the number of epochs before each round of central averaging, $E=1$, each central epoch is equivalent to one epoch in the traditional approach. For the number of subsets, $N=3$, the loss curve reaches its minimum, $0.462$, at around $21$ epochs and then it keeps increasing and reaches0.640.640.640.64 by the end of 100 epochs. When we further divided the training data into more subsets, setting $N=9$, the loss curve reaches its minimum,$0.47$, at around 46 epochs and then it keeps increasing and reaches $0.585$ by the end of 100 epochs. So, clearly, the convergence becomes slower, and the abrupt increase of validation loss is minimized as we divide the training set into more and more subsets, which indicates reduced overfitting.
Also, D2C manages to maintain a lower validation loss in general in all cases, including MNIST as we can see from Figure 8. The minimum value of loss that each of the cases discussed above also tells a similar story. The more subsets we divide the training set into, the less the validation loss we can observe from the curves. When a model maintains a lower validation loss during training, it indicates that the model is able to generalize well to unseen data. Theoretically, overfitting occurs when a model becomes overly complex relative to the data it is learning, allowing it to capture not only meaningful patterns but also specific noise or minor irregularities unique to the training set. This focus on noise causes the model to perform poorly on new data, which is reflected in a higher validation loss or an accelerated rise in validation loss. In contrast, a model that keeps validation loss lower while training likely has a structure that is well-suited to the data’s true underlying patterns without becoming excessively sensitive to its unique characteristics. By maintaining a lower validation loss, the model demonstrates an ability to generalize—learning patterns that extend beyond the training set and apply well to validation data. This suggests that the model has achieved a balance, where it is complex enough to capture essential relationships within the data but not so complex that it becomes reliant on details that do not generalize. Consequently, a model that sustains a lower validation loss is theoretically overfitting less, as it shows a stronger alignment with the data’s general structure rather than its idiosyncrasies. So, we can conclude that our proposed method manages to bring down the validation loss during training compared to the traditional method, which is a definitive sign of better generalization and reduced overfitting. We can further conclude that this reduction in validation loss becomes more apparent as we increase $N$, the number of subsets. The most clear illustration of this phenomenon is seen in the case of FER2013 (Figure 9). The Facial Expression Recognition task using the FER2013 dataset is prone to overfitting to a great extent. Dividing the training set into an increasing number of subsets evidently minimizes that overfitting. So, dividing the training set into more and more subsets helps reduce overfitting, an observation that is consistent with what we saw in Eq. (13). The more the number of subsets ($k$ in that equation), the less the variance, s/k, for the best-case scenario. Hence, with the decrease in variance comes the decrease in overfitting. Also, as we saw before, the true error, represented by validation loss in this case, is minimized using D2C, which proves our hypothesis and shows that D2C is indeed effective in treating the problem of overfitting.
Figure 10: Validation accuracy curves for the traditional method & the D2C method for different numbers of subsets, N, on Fashion MNIST.
However, the loss curve isn’t the only important factor during the training process. The relation between log loss and accuracy may not be so linear. That is why it is important to observe the validation accuracy too. As we can see from Figure 10, for all the different values of the number of subsets, $N$, even though the loss reaches its minimum much earlier, the accuracy shows an increasing trend (even if slightly) almost throughout the entire training process, right up until the 60th epoch. This shows why continuing training for a significant period might be necessary even if the validation loss starts increasing. This also shows why early stopping based on loss curves may not be a good idea in both cases: the traditional approach and the D2C method. This is something we observed in the cases of all these datasets. Also, we can notice a visible improvement in terms of validation accuracy in Figure 10. At the very least, dividing the training set into multiple subsets ensures that the maximum accuracy is attained while keeping the log loss relatively in control, which characterizes a more confident and robust model with better generalization ability.
There is another interesting aspect of the validation loss curve here. As we can see from Figure 8, the validation loss curve generated using the traditional approach deviates much more than the ones generated using the D2C method. A curve that does not fluctuate much after converging suggests that the model has reached a stable state. This stability is often a sign of robustness, meaning the model is not overly sensitive to slight variations in the data or training process. Dividing the dataset into more and more subsets results in smaller and smaller deviations in the validation loss curve. This was evident in previous loss curves too, to a lesser extent. This proves that dividing the training set into multiple subsets not only improves accuracy and loss but also brings stability and consistency to the loss/accuracy curves.
Let us also take a look at the validation loss curves generated using FER2013 while keeping the number of subsets, $N$, constant and varying the number of local epochs before each round of central averaging, $E$. We varied $E$ between $1$ to $3$ and Kept $N$ constant at $2$ and $3$ separately.
Figure 11: Validation loss curves for different numbers of epochs before each round of central averaging, while keeping the number of subsets, N = 2 & 3.
We can see from Figure 11 that when we vary the number of epochs before each round of central averaging, $E$, there is no clear indication of any kind of change in the trend or the rate of change in the validation losses as the training progresses. This is unlike when we varied the number of subsets, $N$ keeping $E$ constant, where we saw that increasing the number of subsets delayed and minimized the increase in validation loss as the training progressed quite clearly. Also, the convergence got slower. But here, for both $N=2$ (shades of blue) and $N=3$ (shades of green), the trends from the validation curves are quite similar for different values of $E$ throughout the training phase. In light of the above observations, we can conclude that varying the number of epochs before each round of central averaging, $E$ does not have any obvious implications on log loss as we saw in the case of varying the number of subsets, $N$. However, an increased number of local epochs before each round of central averaging means more scope for learning from each of the subsets before averaging, sometimes even overfitting on a particular subset a bit more. We will observe the effect of varying the number of epochs further in the later subsections when we observe the final accuracy and log loss in each case. The implication of varying $E$ is likely to be different on different datasets from different domains as we will see in the case of the number of subsets too.
Audio Classification: Overfitting is a common problem in audio classification tasks. But it is more apparent when the model is trained on a particular dataset, and then tested on data from a different source. One way of tackling this issue is by combining datasets to increase the diversity in both training and test/validation data. We combined the datasets TESS [19], CREMA-D [20], and RAVDESS [21] for this task. Let us take a look at the validation loss and accuracy curves generated using this Combined Audio dataset.
Figure 12: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on the Combined Audio dataset.
As we can see from Figure 12, there is overfitting to some extent in all cases, which causes the validation loss to creep up after a few epochs but the rate at which this phenomenon takes place is much less when the training set is divided into multiple subsets. Once again, the rate of increase in validation loss goes down as we increase the number of subsets. Even though we converted the 1D time series audio files to 2D Log-Mel spectrograms to form something like an image, the classification based on these Mel-specs differs greatly from traditional image classification. So, the overfitting-resistant nature of models using our proposed method being apparent in this case too is a very promising sign. However, the small size of this dataset may not be ideal in the context of applying D2C. We will analyze it in detail in the next subsection based on several performance metrics.
Text Classification: Text classification is an important case study when it comes to testing the Divide2Conquer method. For both image classification and audio classification, we used 2D data and CNN to carry out the tasks. But text classification deals with 1D sequences, and as we mentioned in the previous section, we will use bi-directional LSTM in this case. We used AG News [22], a benchmark dataset to test our method primarily in the NLP domain. Let us observe the validation loss and accuracy curves generated using the AG News dataset first.
Figure 13: Validation loss curves for the traditional & the D2C method for different numbers of subsets on AG News.
Figure 14: Validation accuracy curves for the traditional & the D2C method for different numbers of subsets on AG News.
The loss and accuracy curves generated for AG News depict by far the clearest picture in terms of the improvements brought by the proposed method over the traditional monolithic approach. The loss curve from Figure 13 shows the trend we observed in almost all other cases: dividing the training data into more and more subsets delays and minimizes the increase in validation loss as the training progresses. The convergence though, slows down a bit in this case too. More importantly, with no subsets, the validation loss shoots up sharply after around 8 epochs, which indicates overfitting and impacts the accuracy too as the training progresses as we can see from Figure 14. However, when we use D2C, the sharp increase in the validation loss is mitigated drastically. Even with only two subsets, the improvement is quite significant. When the number of subsets is increased further to more than two, the loss curve becomes quite stable, almost completely eliminating the increasing trend and showing consistent improvement as the training progresses. The accuracy curve (Figure 14) also shows clear indications that D2C improves over the traditional approach. With the traditional approach, the accuracy too starts deteriorating after around the 14th epoch, indicating severe overfitting, an issue that no longer exists when we use D2C for any value of $N$ or $E$.
The experiments on all these datasets from 3 different datasets prove that using the D2C method can indeed address overfitting and control the abrupt increase of true error in the form of maintaining a lower validation loss during training. This also shows that D2C works well in different domains and a diverse set of tasks.
VI-C Accuracy, F1 Score, AUC-ROC, and Log Loss
We evaluated our proposed D2C on test sets using Accuracy, F1 Score, Log Loss, and AUC-ROC to capture a comprehensive view of model performance in terms of correctness, class balance, prediction confidence, and discrimination ability. Reporting these final metrics on test sets was essential to confirm that improvements generalized well beyond the training and validation data, verifying true effectiveness in unseen scenarios. We used the average from multiple repetitive runs (to test for repeatability) for each of these metrics. We used the provided test sets with each of the datasets for proper comparisons. In the tables, the top row represents the traditional centralized approach and the rest of the rows represent different variations of the D2C approach. The bold entities represent the best performance in the tables.
Image Classification: Let’s take a look at the performance metrics from the benchmark image classification datasets we primarily used using different settings (the traditional approach and the D2C method with different values of $N$ and $E$).
TABLE II: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on CIFAR-10.
Model Settings Accuracy F1-Score Log Loss AUC-ROC
No Subsets 0.8596 0.8591 0.7346 0.98697
N=2, E=1 0.8611 0.8611 0.7042 0.98693
N=3, E=1 0.8613 0.8609 0.6682 0.98704
N=4, E=1 0.8586 0.8583 0.6516 0.98690
N=5, E=1 0.8552 0.8551 0.6389 0.98628
N=7, E=1 0.8539 0.8535 0.6198 0.98640
N=9, E=1 0.8531 0.8525 0.6068 0.98605
TABLE III: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on Fashion MNIST.
Model Settings Accuracy F1-Score Log Loss AUC-ROC
No Subsets 0.9314 0.9315 0.3467 0.99559
N=2, E=1 0.9323 0.9319 0.3460 0.99549
N=3, E=1 0.9330 0.9329 0.3179 0.99545
N=3, E=2 0.9319 0.9316 0.3303 0.99552
N=3, E=3 0.9337 0.9337 0.3225 0.99590
N=5, E=1 0.9322 0.9321 0.2692 0.99589
N=7, E=1 0.9338 0.9334 0.2467 0.99584
N=7, E=2 0.9305 0.9304 0.2575 0.99600
N=9, E=1 0.9323 0.9323 0.2341 0.99600
TABLE IV: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on MNIST.
Model Settings Accuracy F1-Score Log Loss AUC-ROC
No Subsets 0.9952 0.9952 0.0258 0.99995
N=2, E=1 0.9957 0.9956 0.0271 0.99995
N=3, E=1 0.9958 0.9958 0.0240 0.99998
N=3, E=2 0.9951 0.9950 0.0250 0.99996
N=5, E=1 0.9957 0.9956 0.0202 0.99998
N=7, E=1 0.9951 0.9950 0.0215 0.99997
N=9, E=1 0.9957 0.9957 0.0179 0.99998
N=11, E=1 0.9959 0.9959 0.0161 0.99999
N=11, E=2 0.9958 0.9958 0.0170 0.99998
As we can see from Table II, Table III, and Table IV, the best accuracy, MCC and AUC-ROC were achieved using the D2C method for all three datasets. For CIFAR-10, the number of subsets, $N$, was $3$ for the best model in terms of accuracy. The number of epochs before each round of central averaging, $E$, was just $1$. This approach showed an improvement with an accuracy of $86.13\%$ over the traditional approach with no subsets, which yielded an accuracy of $85.96\%$. Initially, the accuracy improved as we increased the number of subsets, but suffered a consistent drop after it was further increased past $3$. As we discussed, dividing the training set into too many subsets may impact training negatively due to a lack of representative samples. The experimental outcomes show further evidence of that in this case.
For Fashion MNIST too, the best accuracy, MCC, and AUC-ROC were achieved using our proposed Divide2Conquer method. The number of subsets, $N$, was $7$ in this case for the best model in terms of accuracy. The number of epochs before each round of central averaging, $E$, was once again just $1$. This approach again showed an improvement with an accuracy of $93.38\%$ over the traditional approach with no subsets, which yielded an accuracy of $93.14\%$. An increasing trend is visible in accuracy as we increase the number of subsets till it reaches the best conditions with $N=7$.
In the case of MNIST too, the best accuracy, MCC, and AUC-ROC were achieved using our proposed method. The number of subsets, $N$, was $11$ in this case for the best model. The number of epochs before each round of central averaging, $E$, was $1$. The improvement of $0.07$ in error rate over the traditional approach is significant for the MNIST dataset. In this case, the least amount of log loss is also achieved with the best model in terms of accuracy.
We can also see that as we increase the number of subsets, $N$, there is a clear trend of decreasing log loss. As we discussed earlier, increasing the number of subsets further minimizes the effect of noise and outliers. The decrease in log loss is a result of that phenomenon. Even though the best model isn’t usually selected based on log loss values, in the case of similar accuracy (for example, the accuracies for $N=3,E=3$ and $N=7,E=1$ in Fashion MNIST dataset), the best model can be selected based on log loss. Log Loss evaluates how well the model predicts probabilities, penalizing confident but incorrect predictions, which is critical for understanding model reliability. A lower log loss means a more robust and confident model, so the model with a lower log loss is preferred when the test accuracy or F1 score is similar.
The comparison of subset numbers in the best-performing models across CIFAR-10, Fashion MNIST, and MNIST provides critical insights. For CIFAR-10, the optimal model utilized only $3$ training subsets, while the best models for Fashion MNIST and MNIST required $7$ and $11$ subsets, respectively. Notably, increasing the subset count beyond $3$ led to declining accuracy for CIFAR-10—a trend not observed with the other datasets. This discrepancy is likely due to the distinct characteristics of each dataset. While all contain $10$ classes, CIFAR-10, a smaller dataset with RGB images and complex class representations, presents a more challenging classification task compared to the grayscale images of Fashion MNIST and MNIST. Consequently, effective training on CIFAR-10 demands larger sample sizes per subset, and when these subsets become too small, performance degrades. In contrast, Fashion MNIST and MNIST, identical in dataset size, differ in task complexity, with Fashion MNIST presenting a slightly more challenging classification task than MNIST. This complexity results in the optimal Fashion MNIST model requiring fewer subsets than MNIST, as each subset’s sample size needs to be sufficient for adequate training.
This analysis suggests a key principle for applying the D2C method effectively: to divide training data into more subsets, datasets must be large enough to allow each subset ample training data. Larger datasets benefit more from divided training, requiring more subsets to achieve optimal performance.
TABLE V: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on FER2013.
Model Settings Accuracy F1-Score Log Loss AUC-ROC
No Subsets 0.6498 0.6420 2.0112 0.89934
N=2, E=1 0.6511 0.6394 1.9101 0.89875
N=2, E=2 0.6567 0.6444 1.8450 0.9032
N=2, E=3 0.6490 0.6389 1.8718 0.8975
N=3, E=1 0.6491 0.6356 1.6540 0.90104
N=3, E=2 0.6538 0.6465 1.6990 0.89870
N=3, E=3 0.6539 0.6428 1.7315 0.9002
N=5, E=1 0.6500 0.6368 1.5157 0.90153
N=7, E=1 0.6521 0.6432 1.4821 0.90148
FER2013, a challenging dataset for facial expression recognition, represents a task that is particularly prone to overfitting. As seen in Table V, our proposed D2C method outperformed the traditional approach and achieved the highest accuracy, MCC, and AUC-ROC, with the best-performing model using only 2 subsets ($N=2$) and 2 training epochs before each central averaging step ($E=2$). D2C yielded a notable accuracy improvement, reaching 65.67% compared to 64.98% in the traditional model without subsets. Unlike previous datasets, where a single epoch ($E=1$) was optimal, FER2013 benefited from 2 epochs, likely due to the dataset’s complexity, requiring additional training time to capture meaningful features before model averaging.
A decreasing trend in log loss was observed with more subsets, though no consistent pattern emerged with varying epochs ($E$). Additionally, due to FER2013’s class imbalance, accuracy alone may not reliably indicate the best model, so, the F1-score, a metric considering class-wise precision, provides a valuable perspective. Interestingly, the optimal F1-score model (3 subsets, $E=2$) differs from the accuracy-optimized model, with D2C achieving an F1 score of 0.6465 versus 0.642 in the baseline model. This F1-score improvement underscores D2C’s effectiveness in addressing overfitting, particularly in class-imbalanced, complex tasks.
Audio Classification: Now let us take a look at the performance metrics from the Combined Audio dataset.
TABLE VI: Performance Metrics with Traditional Approach vs. Different Settings Using D2C on the Combined Audio Dataset.
Model Settings Accuracy F1-Score Log Loss AUC-ROC
No Subset 0.6953 0.6870 1.3419 0.96570
N=2, E=1 0.6942 0.6879 1.0811 0.96553
N=2, E=2 0.6908 0.6819 1.1636 0.9632
N=3, E=1 0.6920 0.6909 1.2032 0.96416
N=3, E=2 0.6874 0.6830 1.0033 0.96438
N=4, E=1 0.6921 0.6804 1.2840 0.96214
We can see from Table VI that in the Combined Audio dataset, the highest accuracy was achieved using the traditional training approach rather than our proposed D2C method. The traditional model attained a 69.53% accuracy, slightly outperforming the D2C approach, which reached 69.42% with $2$ subsets. Accuracy declined further as the number of subsets increased, likely due to the smaller dataset size (11,632 samples), limiting effective data availability per subset. As a result, dividing the dataset into smaller portions likely impacted performance by reducing the sample count for each class, a significant challenge given the dataset’s complexity and diversity, as it combines samples from multiple sources.
However, despite marginally lower accuracy, the D2C method showed reduced log loss, indicating a trade-off between accuracy and confidence in predictions. In cases of comparable accuracy, lower log loss can be preferable as it indicates more calibrated predictions. Moreover, given the dataset’s class imbalance, the F1-score is a more reliable metric here. The D2C approach yielded a slight improvement in the F1-score (0.6909 vs. 0.687), indicating better generalization across classes. This suggests that D2C supports robust classification with fewer, more distinctive errors, especially in class-imbalanced, complex tasks like this one. In summary, while D2C did not maximize accuracy, it achieved comparable F1 performance with lower log loss, resulting in a model that generalizes well and offers more confident predictions in challenging, class-imbalanced datasets. However, dividing the dataset into too many subsets does worsen the performance of the models in this case.
This analysis leads us to the conclusion that to apply the Divide2Conquer method effectively, the dataset needs to be sufficiently large so that each of the subsets can have enough training data. It is however noteworthy that even with a comparatively smaller dataset, our proposed method still shows some performance improvement in some metrics.
Text Classification: Finally, let us take a look at the performance metrics from the AG News dataset.
TABLE VII: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on AG News.
Model Settings Accuracy F1-Score Log Loss AUC-ROC
No Subset 0.9243 0.9243 0.2611 0.98764
N=2, E=1 0.9291 0.9291 0.2556 0.98873
N=3, E=1 0.9288 0.9288 0.2520 0.98872
N=3, E=2 0.9275 0.9275 0.2567 0.98852
N=5, E=1 0.9293 0.9292 0.2412 0.98881
N=5, E=2 0.9295 0.9296 0.2451 0.98875
N=7, E=1 0.9271 0.9271 0.2353 0.98916
We can observe from Table VII that, once again, the highest accuracy and F1-score were achieved with our proposed D2C method. The number of subsets, $N$, was $5$ in this case for the best model. The number of epochs before each round of central averaging, $E$, was $2$. The D2C method showed a significant improvement with an accuracy of $92.95\%$ over the traditional approach with no subsets, which yielded an accuracy of $92.43\%$. Using the D2C method decreased the error by more than $0.5\%$ in this case. Also, as we can see from Table VII, the trend of a decreasing log loss as we increase the number of subsets is visible here too. The best method in terms of AUC-ROC also happens to be the model with the highest number of subsets (7). It also seems that the log loss increases a bit when we increase the number of epochs, $E$.
The effect of dividing the training set into multiple subsets is the most apparent in this dataset compared to all the previous ones. AG News, the largest dataset in our study, presents an ideal case for D2C, as each class contains 30,000 samples. This abundance of data per class allows for substantial division without sacrificing subset size, which maintains the model’s training stability while leveraging the regularizing effect of partitioning and averaging. Consequently, every variation of D2C (across all values of $N$ and $E$) demonstrated accuracy gains over the traditional approach, underscoring D2C’s suitability for mitigating overfitting in text classification tasks.
In summary, D2C excels with large datasets like AG News, where ample samples per class allow the model to benefit from both training on robust subsets and reducing overfitting through partitioning and averaging. This validates D2C’s potential for significant impact in Big Data applications, where traditional models often struggle to fully exploit available data.
VI-D Augmentation in D2C
Our experiments show that dividing training data into multiple subsets and applying central averaging results in less overfitting and an improvement in performance metrics like accuracy, F1 score, etc. However, we also saw that the training subsets need to be sufficiently large so that each of the subsets can provide enough training diversity to the edge models, otherwise, the performance suffers a drop. When the dataset is large, the proposed method works very well, but when the dataset is smaller, this may cause issues in achieving optimum performance.
One way to counter these issues is data augmentation. This is a widely used technique to expand a dataset and prevent overfitting. To minimize the effect of outliers and noise and to prevent data leakage among the subsets, augmentation must take place after dividing the dataset into multiple subsets. The augmentation is then applied to each of the subsets. That way, the augmented samples from one subset won’t get mixed with another subset. This approach ensures the separation of data in each of the subsets as well as the expansion of training data, providing enough samples for all the subsets, thus improving the overall performance. We tested this augmentation approach on the CIFAR-10 dataset for $3$, $5$ $7$ subsets. The augmentation methods we used were rotation, horizontal flip, and shifting along both axes. Let us take a look at the accuracies we achieved in each case and compare them with the accuracies we got before without applying augmentation.
Figure 15: Test accuracies for different numbers of subsets, N, using the D2C method on the CIFAR-10 dataset with and without Augmentation.
As we can see from Figure 15, the accuracy improved significantly in each case after we applied data augmentation to the subsets of the training data. The accuracy increased from $86.13\%$ to $88.73\%$ for $3$ subsets. In the case of $5$ subsets, the improvement was even more, reaching $88.12\%$ from just $85.52\%$. In the case of $7$ subsets too, the accuracy increased from $85.39\%$ to $87.58\%$. This shows that data augmentation techniques can be successfully applied when using the D2C method, and they can effectively treat the problems that arise when the size of the subsets becomes too small.
VII Limitations
We have shown that our proposed D2C method treats overfitting and shows an improvement over the traditional approach in multiple domains and datasets. However, all these come at a cost. Dividing the training set into multiple subsets and training them all separately means repeating the whole training process $N$ times ($N$ = number of subsets), albeit with less data. This results in an increased training time. Let us take a look at the following Figure 16 for MNIST.
Figure 16: Time elapsed per epoch for the traditional approach and different numbers of subsets, N, using the D2C method on the MNIST dataset.
The training was done using a V100 GPU in each of these cases for fair comparisons. As we can see from Figure 16, the time needed to train the models is the least in the traditional approach, as expected. As we increase the number of subsets, $N$, the time needed to complete each epoch of training increases. However, it is important to note that the increase in training time is not abrupt. It doesn’t increase proportionally with $N$. For example, the time taken to train for one epoch was $12.5$ seconds in the case of the traditional implementation. When the number of subsets, $N$, was $11$, it took $40$ seconds to train for one global epoch. The training time increased significantly, but it was nowhere near $11$ times the time taken in the traditional approach. So, even though the training time increases, it doesn’t shoot up abruptly. It is important because such an abrupt increase in training time could mean a big issue when it comes to applying the Divide2Conquer method.
Lastly, it’s important to note that even if the training time is more when using the D2C method, it doesn’t make any difference at all in the post-training deployment stage, as the edge models (like the one used in the traditional approach) and the central model have the same number of weights, architecture, and computational complexity. So, clearly, even if the Training time increases while using the Divide2Conquer method, the Inference time remains exactly the same as the traditional approach.
VIII Conclusion
Our study demonstrates that dividing training data into multiple subsets and averaging the weights iteratively can treat overfitting and produce significant performance gains through stability and better generalization. In this paper, we applied our proposed Divide2Conquer method and tested it on multiple datasets from different domains. This method outperformed the traditional approach in all the datasets in terms of key metrics and showed significantly better resistance against overfitting. The D2C method also showed a consistent ability to mitigate the increasing trend of validation loss, and to produce smoother decision boundaries. These results prove our hypothesis of this method being able to help generalize better and resist overfitting to be true. Moreover, the Divide2Conquer method brings even more improvement when the dataset is larger, outlining its promise in Big Data scenarios, especially since D2C method can be applied on top of other generalization techniques. Thus, the proposed D2C method opens the door for future research and experiments using this in several domains.
Acknowledgments
This work was supported by the Institute for Advanced Research (IAR), United International University, under Grant UIU-IAR-02-2022-SE-06. The authors sincerely appreciate the funding and resources provided, which were essential to the successful execution of this research.
Appendix A Related Proofs and Derivations
Derivation of $Var(X)=\mathop{{}\mathbb{E}}(X^{2})-(\mathop{{}\mathbb{E}}(X))^{2}$ : Let the random variable $X$ denote the estimate of $x$. The variance of estimating this random variable is defined as the average deviation of $X$ from the expected value of our estimate, $\mathop{{}\mathbb{E}}(X)$, where the deviation is squared for symmetry of difference. So, the variance can be formulated as:
$Var(X)$ $=\mathop{{}\mathbb{E}}((X-\mathop{{}\mathbb{E}}(X))^{2})$
$=\mathop{{}\mathbb{E}}(X^{2}+(\mathop{{}\mathbb{E}}(X))^{2}-2X\mathop{{}\mathbb{E}}(X))$
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\mathop{{}\mathbb{E}}(X^{2})+(\mathop{{}\mathbb{E}}(X))^{2}-2\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(\mathop{{}\mathbb{E}}(X))$
$=\mathop{{}\mathbb{E}}(X^{2})-(\mathop{{}\mathbb{E}}(X))^{2}$
Here, $(a)$ is due to the linear nature of the expectation operator, and $\mathop{{}\mathbb{E}}(\mathop{{}\mathbb{E}}(X))=\mathop{{}\mathbb{E}}(X)$, because $\mathop{{}\mathbb{E}}(X)$ is not a random variable.
Derivation of $CoV(XY)=\mathop{{}\mathbb{E}}(XY)-\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$ : Let $X$ and $Y$ be estimations of two random variables $x$ and $y$ respectively. The covariance of $X$ and $Y$ is defined as the expected value (or mean) of the product of their deviations from their individual expected values[35]. Hence,
$Cov(X,Y)$ $=\mathop{{}\mathbb{E}}((X-\mathop{{}\mathbb{E}}(X)).(Y-\mathop{{}\mathbb{E}}(Y)))$
$=\mathop{{}\mathbb{E}}(XY-X\mathop{{}\mathbb{E}}(Y)-Y\mathop{{}\mathbb{E}}(X)+\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y))$
=$\stackrel{{\scriptstyle\text{(b)}}}{{=}}\mathop{{}\mathbb{E}}(XY)-\mathop{{}\mathbb{E}}(X\mathop{{}\mathbb{E}}(Y))$
$\qquad-\mathop{{}\mathbb{E}}(Y\mathop{{}\mathbb{E}}(X))+\mathop{{}\mathbb{E}}(\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y))$
=$\stackrel{{\scriptstyle\text{(c)}}}{{=}}\mathop{{}\mathbb{E}}(XY)-\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$
$\qquad-\mathop{{}\mathbb{E}}(Y)\mathop{{}\mathbb{E}}(X)+\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$
$=\mathop{{}\mathbb{E}}(XY)-\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$
[…]
=$\stackrel{{\scriptstyle\text{(d)}}}{{=}}\int yf(y)\mathop{{}\mathbb{E}}(x)dy$
[…]
$=\mathop{{}\mathbb{E}}(X)E(Y)$
[…]
Derivation of $Var(\sum_{i=1}^{k}X_{i})=\sum_{i=1}^{k}Var(X_{i})+\sum_{1\leq i\neq j\leq k}Cov(X_{i},X_{j})$:
Let us consider the random variables $X$ and $Y$ and arbitrary constants $a,b$ Using LABEL:proof1 we can write,
$Var(aX+bY)$ $=\mathop{{}\mathbb{E}}((aX+bY)^{2})-(\mathop{{}\mathbb{E}}(aX+bY))^{2}$
$=(a^{2}\mathop{{}\mathbb{E}}(X^{2})+b^{2}\mathop{{}\mathbb{E}}(Y^{2})+2ab\mathop{{}\mathbb{E}}(XY))$
$\qquad-(a\mathop{{}\mathbb{E}}(x)+b\mathop{{}\mathbb{E}}(Y))^{2}$
$=(a^{2}\mathop{{}\mathbb{E}}(X^{2})+b^{2}\mathop{{}\mathbb{E}}(Y^{2})+2ab\mathop{{}\mathbb{E}}(XY))$
$\qquad-a^{2}(\mathop{{}\mathbb{E}}(X))^{2}-b^{2}(\mathop{{}\mathbb{E}}(Y))^{2}$
$\qquad-2ab\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$
=$\stackrel{{\scriptstyle\text{(e)}}}{{=}}a^{2}Var(X)+b^{2}Var(Y)+2abCov(X,Y)$
Here, $(e)$ follows directly from LABEL:proof1 and LABEL:proof2. So Using LABEL:proof3 and LABEL:proof4 and generalizing for $k$ random variables we can say:
$Var(\sum_{i=1}^{k}X_{i})$ $=\sum_{i=1}^{k}Var(X_{i})+\sum_{1\leq i\neq j\leq k}Cov(X_{i},X_{j})$ (18)

### v2

@@
-An Extensive Study on D2C: Overfitting Remediation in Deep Learning Using a Decentralized Approach
-Abstract
-Overfitting remains a significant challenge in deep learning, often arising from data outliers, noise, and limited training data.
-To address this, we propose Divide2Conquer (D2C), a novel technique to mitigate overfitting.
-D2C partitions the training data into multiple subsets and trains identical models independently on each subset.
-To balance model generalization and subset-specific learning, the model parameters are periodically aggregated and averaged during training.
-This process enables the learning of robust patterns while minimizing the influence of outliers and noise.
-Empirical evaluations on benchmark datasets across diverse deep-learning tasks demonstrate that D2C significantly enhances generalization performance, particularly with larger datasets.
-Our analysis includes evaluations of decision boundaries, loss curves, and other performance metrics, highlighting D2C’s effectiveness both as a standalone technique and in combination with other overfitting reduction methods.
-We further provide a rigorous mathematical justification for D2C’s underlying principles and examine its applicability across multiple domains.
-Finally, we explore the trade-offs associated with D2C and propose strategies to address them, offering a holistic view of its strengths and limitations.
-This study establishes D2C as a versatile and effective approach to combating overfitting in deep learning.
-Our codes are publicly available at: https://github.com/Saiful185/Divide2Conquer.
-Index Terms:
-Classification, Deep Learning, Hyperparameter, Overfitting, Regularization, Variance.
-I Introduction
-Deep Learning models often do well on the data they train on, but the performance drops massively on unseen data from different distributions.
-This phenomenon is known as Overfitting.
-Many methods have been proposed to reduce overfitting over the years [10].
-Early Stopping [27] is one of the most intuitive methods for reducing overfitting.
-However, it limits the model’s learning potential [28].
-Network reduction reduces variance by simplifying the model, but it restricts learning of complex features [29].
-Data augmentation is a common approach to mitigate overfitting in deep learning applications [32].
-However, choosing suitable augmentation techniques for specific datasets can be challenging, and obtaining additional training data often requires substantial effort.
-Regularization is a widely used technique to reduce overfitting by penalizing model dependency on training data.
-Dropout, the most relevant regularization method in deep learning, reduces overfitting by randomly dropping connections between neural network layers [4].
-However, even dropout can sometimes degrade the performance of a model due to internal covariance shifts [18].
-Eventually, despite employing these techniques, the model still has access to the entire training data, and the robust neural network structures find their ways to fit on the training set a bit too much.
-This is where our motivation comes from.
-It can be a good idea to not let the neural network train on the entire dataset in the first place.
-Intuitively, we should combine multiple models that train on different portions of the data so that the model can not familiarize itself with all of it at once, and a consensus can be achieved by focusing on the general patterns observed in different portions of the data.
-We propose a novel method, Divide2Conquer (D2C), which proposes dividing the training data into multiple subsets and training a model(all having the same architecture & hyperparameters) on each subset.
-A weighted averaging of all the parameter weights from the instance model is performed after a certain number of epochs, and the averaged weights are shared back to each of the models (we denote them as edge models in this study).
-Then, the whole process is repeated for several global epochs.
-D2C is loosely inspired by Federated Optimization[2], which involves training identical models on local devices (using the data available on those devices) and periodically aggregating and averaging their trained parameters on a central server.
-This process ensures safe access to data and preserves privacy, as it only shares the weights, not the data itself.
-However, unlike D2C, federated learning is not inherently designed to enhance model generalization.
-The D2C method also has similarities with Bagging[23] in the form of using subsets of data.
-However, in bagging (Bootstrap Aggregating), the data is not completely separated for different models as we proposed.
-Instead, the technique involves creating multiple subsets of the original dataset by randomly sampling with replacement.
-This means that some datapoints may be repeated in a subset while others may be excluded altogether.
-Also, in bagging, the aggregation is done by combining the output probabilities, not the model weights.
-Our experiments encompass multiple datasets across different domains, and each case suggests that employing the D2C method can reduce overfitting significantly.
-The primary comparison to evaluate the method was between the performances of the base Neural Network architecture used for the Edge Models using the entire training data and the performance of our models using the Divide2Conquer method.
+DUA-D2C: Dynamic Uncertainty Aware Method for Overfitting Remediation in Deep Learning
+Abstract.
+Background: Overfitting remains a significant challenge in deep learning, often arising from data outliers, noise, and limited training data.
+To address this, we previously proposed the Divide2Conquer (D2C) method, which partitions training data into multiple subsets and trains identical models independently on each.
+This strategy enables learning more consistent patterns while minimizing the influence of individual outliers and noise.
+Objectives: D2C’s standard aggregation typically treats all subset models equally or based on fixed heuristics (like data size), potentially underutilizing information about their varying generalization capabilities.
+Building upon this foundation, we introduce Dynamic Uncertainty-Aware Divide2Conquer (DUA-D2C), an advanced technique that refines the aggregation process.
+Methods: DUA-D2C dynamically weights the contributions of subset models based on their performance on a shared validation set, considering both accuracy and prediction uncertainty.
+This intelligent aggregation allows the central model to preferentially learn from subsets yielding more generalizable and confident edge models, thereby more effectively combating overfitting.
+Results: In this work, we provide a rigorous theoretical justification for DUA-D2C’s efficacy.
+Empirical evaluations on benchmark datasets demonstrate that DUA-D2C significantly improves generalization, even when applied on top of other regularization methods.
+Our analysis includes evaluations of decision boundaries, loss curves, and other performance metrics, highlighting the effectiveness of DUA-D2C.
+Conclusions: This study demonstrates that DUA-D2C improves generalization performance even when applied on top of other methods, establishing it as a theoretically grounded and effective approach to combating overfitting in deep learning.
+copyright: ccdoi: 10.1613/jair.1.xxxxxpublicationmonth: 9journalyear: 2025
+1. Introduction
+Deep Learning models often achieve high performance on their training data, yet this performance can drop significantly on unseen data from different distributions, a phenomenon known as overfitting.
+Numerous methods have been proposed over the years to mitigate overfitting (ying2019overview, 30).
+Early Stopping (morgan1989generalization, 21), while intuitive, may limit a model’s learning potential (prechelt2002early, 24).
+Network reduction simplifies models to reduce variance but can restrict the learning of complex features (furnkranz1997pruning, 7).
+Data augmentation is a common approach (sun2014deep, 28), but selecting suitable techniques can be challenging, and acquiring additional data often requires substantial effort.
+Regularization, particularly methods like Dropout (srivastava2014dropout, 27) which randomly deactivates connections, is widely used.
+However, even Dropout can sometimes degrade performance due to issues like internal covariance shifts (li2019understanding, 17).
+A common thread is that despite these techniques, the model ultimately trains on the entire dataset, allowing robust neural networks to potentially over-adapt to the training set.
+This observation motivates our work.
+We posit that preventing the neural network from training on the entire dataset simultaneously can be beneficial.
+Instead, by combining multiple models trained on different portions of the data, a consensus can be achieved, focusing on general patterns rather than allowing any single model to become overly familiar with the complete dataset.
+To this end, we previously proposed the Divide2Conquer (D2C) method (siddiqui2024d2c, 1).
+D2C partitions the training data into multiple subsets and trains a model (all instances sharing the same architecture and hyperparameters) on each subset.
+After a certain number of local training epochs, the parameters from these edge models (models trained on subsets) are aggregated through weighted averaging, and the resulting averaged weights are disseminated back to each edge model.
+This cycle is repeated for several global epochs.
+In this study, apart from proposing DUA-D2C, we provide a rigorous mathematical justification for the hypothesis that D2C mitigates overfitting by demonstrating its capacity to reduce model variance.
+While D2C provides a robust framework for improving generalization, its standard aggregation mechanism typically treats all edge models equally based on their data subset size.
+To further refine this process and enhance model generalization, we introduce an extension: Dynamic Uncertainty-Aware Divide2Conquer (DUA-D2C).
+DUA-D2C enhances the aggregation step by dynamically weighting the contributions of edge models.
+These weights are determined by evaluating each edge model’s performance on a shared validation set, considering both its predictive accuracy and its prediction uncertainty (e.g., measured via entropy).
+This intelligent, adaptive weighting scheme allows the central model to preferentially learn from edge models that demonstrate superior generalization and confidence, thereby more effectively minimizing the influence of outliers, noise, and subset-specific overfitting.
+The D2C approach was inspired by Federated Optimization (konevcny2015federated, 13), which trains models on local devices and aggregates parameters centrally, primarily for privacy.
+However, unlike D2C and DUA-D2C, federated averaging is not inherently designed as an overfitting reduction technique.
+D2C also shares similarities with Bagging (ghojogh2019theory, 8) in its use of data subsets.
+However, Bagging typically involves sampling with replacement, meaning data is not strictly separated, and aggregation is performed on output probabilities rather than model weights.
+Our experiments encompass multiple datasets across different domains.
+Each case suggests that employing the D2C framework, particularly with the DUA-D2C enhancement, can significantly reduce overfitting.
+The primary comparison to evaluate the methods was between the performance of the base neural network architecture (used for the Edge Models) trained on the entire dataset and the performance of our models utilizing the D2C and DUA-D2C approaches.
 We summarize our contributions through this study below:
-  • We introduce a new method, D2C, that helps in reducing overfitting significantly while being conceptually simple and easy to implement.
-  • We present a mathematical justification for the hypothesis that D2C mitigates overfitting by rigorously demonstrating its capacity to reduce model variance.
-This formal analysis underscores D2C’s potential to improve model robustness across diverse datasets.
-  • D2C can be applied on top of other data augmentation and regularization techniques, and our experiments across various datasets and tasks show that this results in a clear improvement in the model’s generalization ability.
-  • We extensively tune the hyperparameters introduced by this method and report the findings, providing important directions for future applications.
-  • We also delve into the trade-offs associated with D2C, including its computational overhead and potential diminishing returns for smaller datasets.
-We explore strategies to mitigate these challenges and provide a comprehensive understanding of D2C’s strengths and weaknesses.
+  • We introduce Dynamic Uncertainty-Aware D2C (DUA-D2C), a novel extension that employs an intelligent, dynamic weighting scheme based on edge model validation performance (accuracy and uncertainty) to further enhance model robustness and generalization.
+  • We present a rigorous mathematical justification for the Divide2Conquer approach and extend it to highlight DUA-D2C’s effectiveness, demonstrating its ability to reduce model variance and thereby improve generalization.
+  • We demonstrate through extensive experiments that DUA-D2C can be applied on top of other data augmentation and regularization techniques, leading to clear improvements in model generalization across various datasets and tasks.
+  • We delve into the trade-offs associated with the DUA-D2C, such as computational overhead and potential diminishing returns for smaller datasets, exploring strategies to mitigate these challenges and offering a comprehensive understanding of their strengths and weaknesses.
+We also tune the hyperparameters introduced by DUA-D2C and report the findings, providing important directions for future applications.
 The paper is organized as follows.
-In section II, we discuss relevant literature.
-We establish the theoretical justification behind our hypothesis in section III.
-In sections IV and V, we discuss our methodology and lay out the experimental specifications.
-In section VI, we evaluate our approach through empirical experiments on multiple datasets and analyze the results.
-We discuss the limitations of D2C in section VII.
-Section VIII contains concluding remarks.
-II Related Works
-Several authors have addressed the issue of overfitting while performing various tasks.
-M.
-Cogswell et al.
-proposed a regularizer called DeCov, which helps reduce overfitting in deep neural networks by Decorrelating Representations[5].
-Dropout [4] was proposed by Srivastava et al.
-back in 2014, and since then, this technique has been extensively used in very complex neural network architectures successfully.
-It was indeed an outstanding contribution, specifically for deep learning-based models.
-Batch Normalization [8] proposed by Ioffe and Sergei primarily focuses on better convergence and somewhat contributes to reducing overfitting.
-J.
-Kolluri and V.K Kotte came up with L¼ regularization to solve the problems faced by L1 and L2 regularization techniques [15].
-Ensembles are also often used to improve generalization.
-The ensemble-based ELVD model [3] managed to outperform the traditional VGGNet and DropoutNet models in terms of reducing overfitting.
-Zehong Zeng et al.
-came up with an ensemble framework that incorporates several techniques to prevent overfitting [9].
-Min-Gu Kim et al.
-also proposed parallel ensemble networks to reduce overfitting in ECG data and prevent the degradation of generalization performance as the training progresses [1].
-We implemented a method based on federated optimization preliminarily for facial expression recognition using FedNet[17].
-This model achieves excellent results in terms of generalization on both CK+ and FER-2013 datasets.
-The K-Fold-Cross-Validation method for evaluation proposed by Korjus K.
-et al.
-[13] has some similarities to our proposed method in the sense that it also creates subsets by partitioning the dataset and makes use of the whole training data.
-However, this method is different from ours since the erroneous samples still make it to the training phase (K-1/K)-th of the time in K-Fold CV.
-The proposed D2C method partitions data and the whole training process is partitioned too, which means none of the subsets is fed with a particular outlier while training bar one, minimizing the impact of these erroneous samples.
-Overfitting remains a prevalent issue in supervised machine learning despite methods like Early Stopping, Network Reduction, Training Set Expansion, Regularization, and Dropout being effective [10].
-Through D2C, we build an approach that focuses on achieving even better generalization and can be implemented on top of other overfitting-reducing techniques.
-III Theoretical Framework
-III-A The Hypothesis
-Overfitting happens when a model fits too well on the data that it trains on.
-It effectively captures and memorizes even the random characteristics from the training data, randomness that would be insignificant in real-world applications.
-The presence of outliers and noisy data is a fundamental reason behind overfitting.
-In this study, we formulate a method to minimize the effect of these outliers and noisy data.
-The contribution of outliers/noise can be minimized by dividing the training data into multiple shards and training each shard separately.
-That way, each data point will only occur once in one of the several data shards.
-The representative samples would be close to each other in the feature space and would be present in each of the data shards more or less uniformly.
-However, the individual outliers/noise would only be able to impact one of the training processes.
-Aggregation and averaging can be done periodically to combine all the models.
-In the best-case scenario, due to averaging, the effect of the individual outliers/noise would be reduced by a factor of N, where N is the number of data shards.
-The key factors behind our hypothesis and considerations regarding this method being able to address overfitting are discussed as follows:
-1.
-Weighted averaging of parameters helps in combining the knowledge learned from different subsets.
-However, extreme parameter updates driven by noise or outliers in individual models are moderated when averaging across multiple models.
-If a model encounters a noisy sample that deviates significantly from the overall data distribution, it might adjust its parameters excessively to fit that sample, leading to overfitting.
-However, if the training data is divided into subsets and trained separately, even if one model’s parameters are influenced by noisy or outlier samples, their impact is diluted when combined with the parameter weights of other models during averaging.
-So, the averaged parameters should reflect a more balanced representation of the underlying data distribution, mitigating the influence of individual noisy or outlier samples.
-2.
-Averaging the parameter weights is particularly helpful in combating overfitting since it has a regularizing effect.
-Averaging weights regularizes the model at the parameter level.
-It moderates extreme weight updates driven by noise or outliers in individual models, leading to more stable and generalizable parameter values.
-This helps prevent overfitting by discouraging models from fitting to the noise or idiosyncrasies present in their training subsets.
-By averaging parameter weights, the central model converges to a shared solution that reflects the collective knowledge learned from different perspectives.
-This consensus learning approach encourages models to learn generalizable patterns and reduces reliance on individual model predictions that are prone to overfitting.
-3.
-Averaging parameters smooths out the decision boundary learned by individual models.
-Extreme or noisy parameter updates that result in sharp or jagged decision boundaries in individual models are likely to be moderated and smoothed when combined through averaging, consequently helping in generalizing well to unseen data and reducing overfitting.
-4.
-The optimal hyperparameters, such as the number of subsets and epochs before each round of central averaging vary by dataset and domain.
-The appropriate subset count depends on dataset size and class distribution.
-Suppose the edge model is complex or the subset count is too high.
-In that case, the overly reduced size and number of samples per class in each subset can result in significant overfitting in edge models, which in turn can degrade overall performance even after central averaging.
+We establish the theoretical justification behind the D2C hypothesis in Section 2.
+In Section 3, we detail the D2C methodology and its DUA-D2C enhancement, followed by the experimental specifications in Section 4.
+In Section 5, we evaluate our approaches through empirical experiments on multiple datasets and analyze the results.
+We discuss the limitations of our methods in Section 6.
+In Section 7, we discuss relevant literature.
+Section 8 contains concluding remarks.
+2. Theoretical Framework
+2.1. The Hypothesis
+Overfitting occurs when a model adapts excessively to its training data, effectively capturing and memorizing random characteristics and noise that lack significance in real-world applications.
+The presence of outliers and noisy data is a fundamental contributor to this phenomenon.
+Our approach aims to minimize the detrimental impact of such data irregularities.
+The core idea is that the influence of outliers and noise can be substantially mitigated by partitioning the training data into multiple shards and training distinct models on each.
+In this setup, each data point is confined to a single shard.
+While representative samples, being closer in the feature space, are likely to be distributed somewhat uniformly across shards, individual outliers or noisy instances will primarily affect only one of the local training processes.
+Through periodic aggregation of these locally trained models, we can consolidate their learned knowledge.
+For the basic Divide2Conquer (D2C) approach, which employs simple averaging, the adverse effect of individual outliers/noise could ideally be reduced by a factor approaching N (the number of data shards).
+Building upon this, our Dynamic Uncertainty-Aware D2C (DUA-D2C) introduces a more sophisticated aggregation mechanism.
+Instead of uniform averaging, DUA-D2C dynamically weights the contribution of each edge model based on its performance (accuracy) and predictive confidence (inverse uncertainty/entropy) on a shared validation set.
+This refinement aims to further amplify the benefits of the Divide2Conquer paradigm.
+The key factors underpinning our hypothesis that the DUA-D2C method addresses overfitting are as follows:
+I.
+In D2C, weighted averaging of parameters combines knowledge from different subsets.
+Extreme parameter updates driven by noise or outliers in individual models are moderated during this averaging.
+If an edge model encounters a noisy sample that deviates significantly, it might adjust its parameters excessively.
+However, this impact is diluted when combined with parameters from other models.
+The DUA-D2C extension enhances this by actively down-weighting models that exhibit poor generalization (low accuracy on validation) or high uncertainty (potentially due to fitting noise in their local shard).
+This ensures that the aggregated parameters more reliably reflect a balanced representation of the underlying data distribution, offering a more robust mitigation against individual noisy or outlier samples.
+II.
+Averaging weights has a regularizing effect at the parameter level.
+D2C’s averaging moderates extreme weight updates, leading to more stable parameter values.
+DUA-D2C refines this by making the regularization adaptive.
+By prioritizing models that are both accurate and confident on unseen validation data, DUA-D2C discourages the central model from adopting parameters heavily influenced by noise or idiosyncrasies from less reliable edge models.
+This intelligent consensus learning encourages the central model to learn more generalizable patterns and reduces reliance on individual model predictions that might be prone to overfitting their local data.
+III.
+Averaging parameters generally smooths out the decision boundaries learned by individual models.
+Sharp or jagged boundaries resulting from noisy updates in individual models are moderated through D2C’s averaging.
+DUA-D2C provides an additional layer of quality control: models that likely developed overly complex or noisy decision boundaries (as indicated by poor validation performance or high uncertainty) contribute less to the central model.
+This leads to a central model with a potentially smoother and more robust decision boundary, better equipped for generalization to unseen data.
+IV.
+The optimal hyperparameters, such as the number of subsets $N$ and local epochs $E$, remain crucial and vary by dataset and domain.
+The subset count depends on dataset size and class distribution.
+If edge models are complex or the subset count is too high, the reduced data per subset can lead to significant overfitting in these edge models.
 This is similar to what happens in federated learning when the client data is minimal, as pointed out by Zhang et al.
-[34].
-When sufficient data exists per class, more subsets can be beneficial, as noise or outliers are diluted by a factor of up to N (number of training subsets).
-In summary, D2C helps to reduce overfitting by mitigating the effects of noise and outliers.
-However, dividing the training set into too many subsets can lead to insufficient data per subset, which can hamper the overall performance.
-III-B Mathematical Intuition
+(zhang2022fed, 34).
+Statistically, the variance of an estimated model $Var(\hat{f}_{j})$ is often inversely related to the amount of data used for its estimation.
+When $n_{j}$ is small, the parameters $\theta_{j}$ learned by $\hat{f}_{j}$ can be highly sensitive to the specific instances in $D_{j}$.
+Small perturbations in these few samples can lead to large changes in the learned $\theta_{j}$, indicating high variance.
+The model might learn spurious correlations present only in that small $D_{j}$ or effectively memorize its samples, rather than capturing the underlying, generalizable patterns from the true data distribution $P(x,y)$.
+This local overfitting manifests as $Var(\hat{f}_{j})$ being large, meaning $\hat{f}_{j}$ would produce widely different predictions if trained on a different small random sample from $P(x,y)$.
+When sufficient data exists per class ($n_{j}$ is reasonably large), more subsets can be beneficial, as noise or outliers are diluted by a factor of up to N (number of training subsets), and each $\hat{f}_{j}$ can learn more stable representations.
+While D2C’s averaging helps, DUA-D2C offers a mechanism to better handle this scenario.
+If some edge models overfit due to limited data in their shard, their likely poor performance or high uncertainty on the validation set will reduce their influence on the central model, offering a degree of resilience against poorly performing edge models.
+Nevertheless, ensuring sufficient data per class in each subset remains important for all edge models to learn meaningful representations.
+The dynamic weighting in DUA-D2C aims to leverage the ”best” of what each edge model learns, even if some are suboptimal.
+In summary, DUA-D2C’s dynamic uncertainty-aware aggregation helps reduce overfitting by mitigating the effects of noise and outliers and by promoting a more robust, generalized central model.
+While dividing the training set into too many subsets can still pose challenges, DUA-D2C provides a more sophisticated mechanism to capitalize on the contributions of potentially variably performing edge models.
+2.2. Mathematical Intuition
 To lay out the underlying intuition behind our hypothesis, we must first define overfitting formally.
 We follow the groundwork laid out by Ghojogh et al.
-[23].
+(ghojogh2019theory, 8).
 Let us consider a Dataset $D$ with training set $T$ and test set $R$ so that,
-$T\cup R=D$ (1)
-$T\cap R=\emptyset$ (2)
+(1) $T\cup R=D$
+(2) $T\cap R=\emptyset$
 Let, $T$ consist of datapoints $\{(x_{i},y_{i})\}_{i=1}^{n}$ and test set $R$ consist the datapoints $\{(x_{i},y_{i})\}_{i=1}^{m}$.
 Also, let $f$ be the function that maps each datapoint in the training set to its corresponding output label(true observations), i.e., $f(x_{i})=f_{i}$, where $f_{i}$ is the true corresponding output label for $x_{i}$.
-This model is unknown and we wish to train a model that replicates $f$ as closely as possible.
+This model is unknown, and we wish to train a model that replicates $f$ as closely as possible.
 However, both the ground truth model and the true labels are unknown to us.
@@
 So, we can say,
-$y_{i}=f_{i}+\epsilon_{i}$ (3)
+(3) $y_{i}=f_{i}+\epsilon_{i}$
 where $\epsilon_{i}$ is some Gaussian noise, $\epsilon\sim\mathcal{N}(0,\sigma^{2})$.
 Since the mean of the distribution is zero we can say that $\mathop{{}\mathbb{E}}(\epsilon_{i})=0$, and $Var(\epsilon_{i})=\sigma^{2}$.
-As the variance of a random variable $x$ can be denoted as $Var(X)=\mathop{{}\mathbb{E}}(x^{2})-[\mathop{{}\mathbb{E}}(x)]^{2}$ (Appendix A), we have,
-$\mathop{{}\mathbb{E}}(\epsilon_{i}^{2})=Var(\epsilon_{i})+[\mathop{{}\mathbb{E}}(\epsilon_{i})]^{2}\Rightarrow\mathop{{}\mathbb{E}}(\epsilon_{i}^{2})=\sigma^{2}+0=\sigma^{2}$ (4)
+As the variance of a random variable $x$ can be denoted as $Var(X)=\mathop{{}\mathbb{E}}(x^{2})-[\mathop{{}\mathbb{E}}(x)]^{2}$ (see Appendix A), we have,
+(4) $\mathop{{}\mathbb{E}}(\epsilon_{i}^{2})=Var(\epsilon_{i})+[\mathop{{}\mathbb{E}}(\epsilon_{i})]^{2}\Rightarrow\mathop{{}\mathbb{E}}(\epsilon_{i}^{2})=\sigma^{2}+0=\sigma^{2}$
 Our assumption is that the input of the training dataset ${\{{x}_{i}}\}_{i=1}^{n}$ and their noise-added outputs ${\{{y}_{i}}\}_{i=1}^{n}$ are available to us, and our goal is to estimate the true model $f$ by a model $\hat{f}$ in order to predict the labels ${\{{y}_{i}}\}_{i=1}^{n}$ for the input data ${\{{x}_{i}}\}_{i=1}^{n}$.
@@
 It is preferable that ${\{\hat{y}_{i}}\}_{i=1}^{n}$ are as close as possible to ${\{{y}_{i}}\}_{i=1}^{n}$.
-Mathematically, we can consider this as a minimization problem looking to minimize some error function such as MSE (Mean squared error).
+Mathematically, we can consider this as a minimization problem looking to minimize some error function, such as MSE (Mean Squared Error).
 The MSE of estimated outputs from the trained model with respect to the dataset outputs are: $\textbf{MSE}=\frac{1}{n}\sum_{i=1}^{n}(\hat{y}_{i}-y_{i})^{2}=\mathop{{}\mathbb{E}}((\hat{y}_{i}-y_{i})^{2})$.
-Let us consider a single arbitrary datapoint $(x_{0},y_{0})$ and the corresponding prediction of the trained model $\hat{f}$ for the input $x_{0}$ is $\hat{y}_{0}$.
+Consider a single arbitrary data point $(x_{0},y_{0})$ and the corresponding prediction of the trained model $\hat{f}$ for the input $x_{0}$ is $\hat{y}_{0}$.
 Ghojogh et al.
-[23] showed that if the sample instance $(x_{0},y_{0})\notin T$ i.e., it belongs to the test set $R$, MSE of the predicted label can be expressed as -
-$\textbf{MSE}=\mathop{{}\mathbb{E}}((\hat{y}_{0}-y_{0})^{2})=\mathop{{}\mathbb{E}}((\hat{y}_{0}-f_{0})^{2})+\sigma^{2}$ (5)
+(ghojogh2019theory, 8) showed that if the sample instance $(x_{0},y_{0})\notin T$ i.e., it belongs to the test set $R$, MSE of the predicted label can be expressed as -
+(5) $\textbf{MSE}=\mathop{{}\mathbb{E}}((\hat{y}_{0}-y_{0})^{2})=\mathop{{}\mathbb{E}}((\hat{y}_{0}-f_{0})^{2})+\sigma^{2}$
 Here we can observe from the first term of the R.H.S.
 of the equation that the MSE of the estimation of output labels with respect to the dataset labels can actually be represented by the MSE of the estimation with respect to the true uncorrupted labels ${\{{f}_{i}}\}_{i=1}^{n}$ plus some effect of the noise that exists in the dataset.
-Taking Monte-Carlo approximation [24] of the expectation terms in Eq.
+Taking Monte-Carlo approximation (ghojogh2020sampling, 9) of the expectation terms in Eq.
 (5), we have:
 $\frac{1}{m}\sum_{i=1}^{m}{(\hat{y}_{i}-y_{i})^{2}}=\frac{1}{m}\sum_{i=1}^{m}{(\hat{y}_{i}-f_{i})^{2}}+\sigma^{2}$
-$\implies\sum_{i=1}^{m}{(\hat{y}_{i}-y_{i})^{2}}=\sum_{i=1}^{m}{(\hat{y}_{i}-f_{i})^{2}}+m\sigma^{2}$ (6)
+(6) $\implies\sum_{i=1}^{m}{(\hat{y}_{i}-y_{i})^{2}}=\sum_{i=1}^{m}{(\hat{y}_{i}-f_{i})^{2}}+m\sigma^{2}$
 Here the term $\sum_{i=1}^{m}{(\hat{y}_{i}-y_{i})^{2}}$ is the total error between the model predictions and the dataset labels which we can call the empirical observed error, $e$.
@@
 But, as Ghojogh et al.
-[23] showed, if $(x_{0},y_{0})\in T$, deriving the M.S.E yields:
-$\textbf{MSE}=\mathop{{}\mathbb{E}}((\hat{y}_{0}-y_{0})^{2})=\mathop{{}\mathbb{E}}((\hat{y}_{0}-f_{0})^{2})+\sigma^{2}-2\sigma^{2}\mathop{{}\mathbb{E}}(\frac{\partial\hat{y}_{0}}{\partial y_{0}})$ (7)
+(ghojogh2019theory, 8) showed, if $(x_{0},y_{0})\in T$, deriving the M.S.E yields:
+(7) $\textbf{MSE}=\mathop{{}\mathbb{E}}((\hat{y}_{0}-y_{0})^{2})=\mathop{{}\mathbb{E}}((\hat{y}_{0}-f_{0})^{2})+\sigma^{2}-2\sigma^{2}\mathop{{}\mathbb{E}}(\frac{\partial\hat{y}_{0}}{\partial y_{0}})$
 Using the same approximation as Eq.(6) on the training set $T$, we get:
 $\frac{1}{n}\sum_{i=1}^{n}{(\hat{y}_{i}-y_{i})^{2}}=\frac{1}{n}\sum_{i=1}^{n}{(\hat{y}_{i}-f_{i})^{2}}+\sigma^{2}-2\sigma^{2}\frac{1}{n}\sum_{i=1}^{n}\frac{\partial\hat{y}_{i}}{\partial y_{i}}$
-$\implies\sum_{i=1}^{n}{(\hat{y}_{i}-y_{i})^{2}}=\sum_{i=1}^{n}{(\hat{y}_{i}-f_{i})^{2}}+n\sigma^{2}-2\sigma^{2}\sum_{i=1}^{n}\frac{\partial\hat{y}_{i}}{\partial y_{i}}$ (8)
+(8) $\implies\sum_{i=1}^{n}{(\hat{y}_{i}-y_{i})^{2}}=\sum_{i=1}^{n}{(\hat{y}_{i}-f_{i})^{2}}+n\sigma^{2}-2\sigma^{2}\sum_{i=1}^{n}\frac{\partial\hat{y}_{i}}{\partial y_{i}}$
 From Eq.(8) we now understand while the model is training the observed empirical error calculated on the training dataset i.e, training error is not a clear representation of the true error because now $e=E+n\sigma^{2}-2\sigma^{2}\sum_{i=1}^{n}\frac{\partial\hat{y}_{0}}{\partial y_{0}}$.
@@
 Figure 2 depicts a visualization of this idea.
-Figure 1: When the corresponding complexity term grows and dominates the expression in Eq.(8) for empirical error, the model starts to overfit as empirical error becomes no longer a true representation of true error.
-Figure 2: A visualization of (a) Underfitting, (b) A good fit, and (c) Overfitting scenarios, illustrating how the model changes if an arbitrary training data point $(x,y)$ is shifted to $(p,q)$ by plotting the trained model before (solid line) and after (dashed line) the shift.
+Figure 1.
+When the corresponding complexity term grows and dominates the expression in Eq.(8) for empirical error, the model starts to overfit as empirical error becomes no longer a true representation of true error.
+Figure 2.
+A visualization of (a) Underfitting, (b) A good fit, and (c) Overfitting scenarios, illustrating how the model changes if an arbitrary training data point $(x,y)$ is shifted to $(p,q)$ by plotting the trained model before (solid line) and after (dashed line) the shift.
 Statistical analysis shows that the distinction between an underfitting model, an overfitting one, and a good fit can be reflected by their bias and variance with respect to the training labels.
@@
 It usually happens when the model is too simple to capture the complexity of the data or has not been trained for long enough to reach convergence.
-On the other hand, we find a model which overfits has high variance and low bias.
-This is because the model becomes overly complex accommodating for all the outliers in the training data and fails to generalize on unseen test data.
-Figure 2(a) and 2(c) show how an underfitting and overfitting model shows low and high variance respectively.
+On the other hand, we find a model that overfits has high variance and low bias.
+This is because the model becomes overly complex, accommodating all the outliers in the training data, and fails to generalize to unseen test data.
+Figure 2(a) and 2(c) show how an underfitting and overfitting model show low and high variance, respectively.
 Figure 2 is a visualization of (a) Underfitting, (b) A good fit, and (c) Overfitting scenarios.
@@
 In (a) the underfitted model the change is very minimal.
-In the (b) Good fit model the change is also relatively little, but in the (c) Overfitted model there is a large shift in the neighboring region of that data point, which reflects the over-dependence on training data.
-To see the impact of D2C we consider the training set $T$ divided into $k$ equal partitions, namely $\lvert T_{1}\rvert=\lvert T_{2}\rvert=...=\lvert T_{k}\rvert=\lvert T\rvert/k$.
+In the (b) Good fit model the change is also relatively little, but in the (c) Overfitted model, there is a large shift in the neighboring region of that data point, which reflects the over-dependence on training data.
+To see the impact of Divide2Conquer we consider the training set $T$ divided into $k$ equal partitions, namely $\lvert T_{1}\rvert=\lvert T_{2}\rvert=...=\lvert T_{k}\rvert=\lvert T\rvert/k$.
 Then, we train the model $\hat{f}_{j}$ using the $j$-th data chunk, where $j\in\{1,..,k\}$.
@@
 For an input vector $\boldsymbol{x}=\{x_{1},...,x_{n}\}$ the prediction of the edge model on the $t$-th epoch can be described as: $\hat{f}_{j}(\boldsymbol{x})=\boldsymbol{W_{j}}^{t}\boldsymbol{x}^{T}+b_{j}$.
-In the case of a single global epoch, we now wish to aggregate the weight matrices creating the central weight matrix $\boldsymbol{\bar{W}}=\{\bar{w}_{1},...,\bar{w}_{n}\}$.
+In the case of a single global epoch, we now wish to aggregate the weight matrices, creating the central weight matrix $\boldsymbol{\bar{W}}=\{\bar{w}_{1},...,\bar{w}_{n}\}$.
 Here, the weight matrix for the $(t+1)$-th epoch is:
-$\boldsymbol{\bar{W}}^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\boldsymbol{W_{j}}^{t}\textrm{ and }b^{t+1}=\frac{1}{k}\sum_{j=1}^{k}{b^{t}_{j}}$ (9)
+(9) $\boldsymbol{\bar{W}}^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\boldsymbol{W_{j}}^{t}\textrm{ and }b^{t+1}=\frac{1}{k}\sum_{j=1}^{k}{b^{t}_{j}}$
 Hence, the output of the central model at $(t+1)$-th epoch:
@@
 $\implies\hat{f}(\boldsymbol{x})^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\left(\boldsymbol{W_{j}}^{t}\boldsymbol{x}^{T}+{b^{t}_{j}}\right)$
-$\implies\hat{f}(\boldsymbol{x})^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\hat{f}(\boldsymbol{x})^{t}_{j}$ (10)
+(10) $\implies\hat{f}(\boldsymbol{x})^{t+1}=\frac{1}{k}\sum_{j=1}^{k}\hat{f}(\boldsymbol{x})^{t}_{j}$
 Let $e_{jt}$ denote the empirical error of the $j$-th model while estimating some arbitrary instance on the $t$-th global epoch.
@@
 Hence, $\mathop{{}\mathbb{E}}(e_{jt})=0$ and $Var(e_{jt})=s$.
-Let us denote the estimation of the product of two trained models using two different data chunks by $c$, i.e, $\mathop{{}\mathbb{E}}(e_{jt}\times e_{lt})=c$.
+Let us denote the estimation of the product of two trained models using two different data chunks by $c$, i.e., $\mathop{{}\mathbb{E}}(e_{jt}\times e_{lt})=c$.
 So, we have:
-$\mathop{{}\mathbb{E}}(e_{jt}^{2})=Var(e_{jt})+[\mathop{{}\mathbb{E}}(e_{jt})]^{2}\Rightarrow\mathop{{}\mathbb{E}}(e_{jt}^{2})=s+0=s$ (11)
+(11) $\mathop{{}\mathbb{E}}(e_{jt}^{2})=Var(e_{jt})+[\mathop{{}\mathbb{E}}(e_{jt})]^{2}\Rightarrow\mathop{{}\mathbb{E}}(e_{jt}^{2})=s+0=s$
 Again for $j,l\in\{1,...,k\}$, where $j\neq l$, we have:
@@
 $\implies Cov(e_{jt},e_{lt})=c-0\times 0=c$
-$\implies Cov(\hat{f}(\boldsymbol{x})^{t}_{j},\hat{f}(\boldsymbol{x})^{t}_{l})=c-0\times 0=c$ (12)
+(12) $\implies Cov(\hat{f}(\boldsymbol{x})^{t}_{j},\hat{f}(\boldsymbol{x})^{t}_{l})=c-0\times 0=c$
 From Eq.(10) and Eq.(12), we can deduce:
-$Var(\hat{f}(\boldsymbol{x})^{t+1})$ $=\frac{1}{k^{2}}Var\left(\sum_{j=1}^{k}\hat{f}(\boldsymbol{x})^{t}_{j}\right)$ (13)
+(13) $Var(\hat{f}(\boldsymbol{x})^{t+1})$ $=\frac{1}{k^{2}}Var\left(\sum_{j=1}^{k}\hat{f}(\boldsymbol{x})^{t}_{j}\right)$
 $=\frac{1}{k^{2}}\sum_{j=1}^{k}Var\left(\hat{f}(\boldsymbol{x})^{t}_{j}\right)+\frac{1}{k^{2}}\sum_{i=1,l=1}^{k}$
@@
 $=\frac{1}{k^{2}}sk+\frac{1}{k^{2}}ck(k-1)=\frac{s+c(k-1)}{k}$
-Relevant proofs related to the derivation of Eq.
-12 and 13 are provided in Appendix A.
 Eq.(13) provides us with an interesting insight.
-If the two edge models are highly correlated i.e., their predictions and empirical error are very close to each other we can assume $s\approx c$.
+If the two edge models are highly correlated, i.e., their predictions and empirical error are very close to each other, we can assume $s\approx c$.
 Hence, the variance of the central model at $(t+1)$-th epoch can be reduced to: $Var(\hat{f}(\boldsymbol{x})^{t+1})\approx(s+s(k-1))/k\approx s$.
@@
 This result is similar to the one proposed by Breiman et al.
-[25] and Bühlmann et al.
-[26] through the idea of ’Bagging’, which is a meta-algorithm aggregating multiple model outputs.
-The analysis clearly shows that our approach does not at the least worsen the problem of overfitting in this scenario but rather improves the central model on each epoch if certain conditions are met.
+(breiman1996bagging, 2) and Bühlmann et al.
+(buhlmann2002analyzing, 3) through the idea of ’Bagging’, which is a meta-algorithm aggregating multiple model outputs.
+The analysis clearly shows that our approach does not at least worsen the problem of overfitting in this scenario, but rather improves the central model on each epoch if certain conditions are met.
 However, in reality, the edge models are neither completely different nor exactly the same due to the relative distribution of the data subsets and their shared model architecture.
@@
 However, it is easy to see that if the training data is contaminated by outliers and random noise, the learned parameters from the different portions of the data are more likely to be different from each other.
-This is because the individual outliers and noise are not highly correlated, hence distributing them among the subsets likely results in significantly different samples in the subsets, making the subsequent models different from each other.
+This is because the individual outliers and noise are not highly correlated; hence, distributing them among the subsets likely results in significantly different samples in the subsets, making the subsequent models different from each other.
 Our results show that repeating the process in every global epoch reduces overfitting in most cases.
-A limitation of this analysis is that deep neural networks are mostly black boxes and their architectures vary tremendously, which makes establishing a generalized mathematical representation of them extremely difficult.
-Hence, our assumption of a simple regression model is not an exact reflection of the actual implementation.
-Despite that, it describes an ideal and simple scenario where the proposed method should improve the model.
-Moreover, it also gave us the insight that this approach can be used in conjunction with other overfitting reduction techniques such as Dropouts(Srivastava et al., 2014) [4].
-If dropouts are stacked on top of our method in every iteration of local training, the neurons are randomly removed with some probability $p$(usually $p=0.5$).
-This introduction of randomness ensures that edge models are different from each other increasing the chances of improvement when the models are aggregated.
-IV Proposed Methodology
-IV-A Creating Training Subsets
-To train multiple Neural Networks parallelly using different subsets of the data, we need to divide the whole training set into multiple shards, and we also need multiple identical models that will be used to train the different subsets.
-These models will be denoted as Edge models in this paper.
-At first, the training set is randomly shuffled and then divided into multiple subsets, maintaining consistent class distribution.
-Each subset is then fed into a separate edge model with identical neural network architecture and trainable parameters to enable parameter averaging.
-IV-B Training & Averaging of Trained Parameters
-Instead of one training loop, the implementation of our method requires two loops: an inner loop for subset training which is nested within an outer loop that performs central averaging.
-A central model with the same architecture as the edge models is first initialized.
-In total, we have one Central Model and $N$ Edge models, all of which share an identical architecture, where $N$ is the number of subsets we have divided the training set into.
-Initially, these identical model objects are built and loaded for each training subset.
-Within the inner loop, each edge model is iteratively trained on its subset for a specified number of local epochs.
-Then, the edge models’ parameters are scaled based on their data fraction and added to a local weight list.
-In the outer loop, these scaled parameters are averaged, updating the central model’s parameters.
-Each edge model’s weights are then reinitialized to match the central model, completing a global training epoch.
-This process repeats over several global epochs, with weighted parameter averaging to account for the data fraction of each subset.
-The whole process is summarized in Figure 3, and the pseudo-code to implement D2C is provided in Algorithm 1.
-Algorithm 1 Divide2Conquer
-  Input: Training dataset $D$, Number of subsets $N$, Local epochs $E$, Global epochs $E_{global}$, Batch size $B$, Learning rate $lr$
-  Output: Central model $M_{c}$ with aggregated weights
-  function Divide2Conquer($D$, $N$, $E$, $E_{global}$, $B$, $lr$)
-        Shuffle the dataset $D$
-        Divide $D$ into $N$ subsets: $\{D_{1},D_{2},\dots,D_{N}\}$, maintaining class distributions
-        Initialize central model $M_{c}$ with parameter vector $\theta_{c}$
-  for $i=1$ to $N$ do
-           Initialize edge model $M_{i}$ with $\theta_{i}=\theta_{c}$
-           Set scaling factor $s_{i}=\frac{|D_{i}|}{|D|}$
-  end for
-  for $g=1$ to $E_{global}$ do
-           Initialize weight accumulator $W_{central}=0$
-     for $i=1$ to $N$ do
-                  Set $\theta_{i}=\theta_{c}$
-        for $e=1$ to $E$ do
-                         Train $M_{i}$ on $D_{i}$ using batch size $B$ and learning rate $lr$
-        end for
-                  Obtain updated weights $\theta_{i}$ from $M_{i}$
-                  Scale weights: $\theta_{i}=s_{i}\cdot\theta_{i}$
-                  Accumulate scaled weights into $W_{central}$: $W_{central}=W_{central}+\theta_{i}$
-     end for
-           Compute averaged central weights:
-$\theta_{c}=\frac{1}{\sum_{i=1}^{N}s_{i}}W_{central}$
-           Update central model $M_{c}$ with new weights $\theta_{c}$
-  end for
-  return central model $M_{c}$
-  end function
-IV-C Global Hyperparameter Tuning
-Tuning the overall model consists of two stages.
-Firstly, the edge model and its hyperparameters should be tuned using a subset/entire training set and the validation set to find a suitable model to train each training subset.
-Next, that base model is utilized to tune the Central model.
-The new global hyperparameters that need tuning, in this case, are the Number of Subsets of the Training Set and the Number of Epochs before each round of Global Averaging.
-Testing for different values for these hyperparameters, the implementation, and the evaluation of our method are done using the best combination.
-The appropriate number of training subsets can be determined first by varying the number of training subsets and observing the performance metrics like test/validation accuracy, F1 score, log loss, and also the validation loss curve.
-The appropriate number of training epochs in each subset before the averaging is done each time is likely to be a small number $(1\sim 3)$ since the edge model can quickly overfit the training subset with a reduced size.
-So, it makes more sense to set a small number of epochs initially and tune the number of training subsets.
-Once the number of subsets is determined, the same process can be repeated by varying the number of training epochs and determining the appropriate value of it.
-Figure 3: D2C at a Glance – Training on each subset, averaging the trained parameters, setting the Central model parameters to these values, and sharing the averaged parameters back to the Edge models to keep the loop running.
-V Experimental Setup
-V-A Datasets
-For image classification tasks, we employed three widely used benchmark datasets: CIFAR-10 [14], MNIST [7], and Fashion MNIST [12].
-CIFAR-10 provides a challenging image classification task with 60,000 32x32 RGB images across 10 categories.
-MNIST offers a simpler task with 70,000 28x28 grayscale images of handwritten digits.
-Fashion MNIST presents a more complex variation with 70,000 28x28 grayscale images of fashion items.
-To address the overfitting challenge in facial expression recognition, we utilized the FER-2013 [16] dataset, which comprises over 35,000 grayscale images of seven basic emotions.
+To provide tractable mathematical intuition for the proposed DUA-D2C/D2C framework, we have adopted a simplified linear regression model.
+This approach is a common and necessary simplification in theoretical analyses within the machine learning literature, utilized in seminal works such as (ghojogh2019theory, 8).
+While this linear model does not fully capture the non-linear complexities of the deep neural networks implemented in our experiments, it serves as a valuable and interpretable analogue.
+It allows us to isolate and illustrate the core statistical principles, such as variance reduction through averaging and the potential benefits of weighted aggregation, that underpin why D2C and DUA-D2C are hypothesized to improve generalization and mitigate overfitting.
+The insights derived from this simplified model provide a foundational understanding, which is then empirically validated through extensive experiments on complex, real-world deep learning architectures.
+Moreover, it also gave us the insight that this approach can be used in conjunction with other overfitting reduction techniques such as Dropouts(Srivastava et al., 2014) (srivastava2014dropout, 27).
+If dropouts are stacked on top of our method in every iteration of local training, the neurons are randomly removed with some probability pp(usually $p=0.5$).
+This introduction of randomness ensures that edge models are different from each other, increasing the chances of improvement when the models are aggregated.
+2.3. Weighted Aggregation
+The DUA-D2C method refines the aggregation step by introducing a dynamic, data-driven weighting scheme for the edge models, rather than a simple uniform average.
+Let $\alpha_{j}$ be the normalized weight assigned to the $j$-th edge model $M_{j}$ (with parameters $\boldsymbol{W_{j}}^{t},b_{j}^{t}$) at global epoch $t$.
+This weight $\alpha_{j}$ is derived from the model’s performance on a validation set, considering its accuracy $a_{j}$ and its inverse prediction uncertainty $u_{j}$ (scaled confidence), such that $\sum_{j=1}^{k}\alpha_{j}=1$ and $\alpha_{j}\geq 0$.
+The aggregation for the central model’s parameters in DUA-D2C at epoch $(t+1)$ becomes:
+(14) $\boldsymbol{\bar{W}}^{t+1}=\sum_{j=1}^{k}\alpha_{j}\boldsymbol{W_{j}}^{t}\quad\text{and}\quad b^{t+1}=\sum_{j=1}^{k}\alpha_{j}{b^{t}_{j}}$
+Consequently, the output of the DUA-D2C central model at epoch $(t+1)$ is:
+$\hat{f}(\boldsymbol{x})^{t+1}$ $=\sum_{j=1}^{k}\alpha_{j}\left(\boldsymbol{W_{j}}^{t}\boldsymbol{x}^{T}+{b^{t}_{j}}\right)$
+(15) $=\sum_{j=1}^{k}\alpha_{j}\hat{f}(\boldsymbol{x})^{t}_{j}$
+This is now a weighted average of the predictions of the edge models. The variance of this weighted sum can be expressed as:
+(16) $Var(\hat{f}(\boldsymbol{x})^{t+1})$ $=Var\left(\sum_{j=1}^{k}\alpha_{j}\hat{f}(\boldsymbol{x})^{t}_{j}\right)$
+$=\sum_{j=1}^{k}\alpha_{j}^{2}Var\left(\hat{f}(\boldsymbol{x})^{t}_{j}\right)+\sum_{i\neq l}\alpha_{i}\alpha_{l}Cov\left(\hat{f}(\boldsymbol{x})^{t}_{i},\hat{f}(\boldsymbol{x})^{t}_{l}\right)$
+$=s\sum_{j=1}^{k}\alpha_{j}^{2}+c\sum_{i\neq l}\alpha_{i}\alpha_{l}$
+assuming, as before, $Var(\hat{f}(\boldsymbol{x})^{t}_{j})=s$ for all $j$, and $Cov(\hat{f}(\boldsymbol{x})^{t}_{i},\hat{f}(\boldsymbol{x})^{t}_{l})=c$ for $i\neq l$.
+The key insight from Eq.
+(16) is that the contribution of each edge model’s variance ($s$) and covariance ($c$) to the central model’s variance is now scaled by the square of its weight ($\alpha_{j}^{2}$) or the product of weights ($\alpha_{i}\alpha_{l}$).
+DUA-D2C aims to assign higher $\alpha_{j}$ values to edge models that generalize better (higher accuracy on a validation set) and are more confident (lower uncertainty).
+  • Prioritizing Lower-Variance Models (Implicitly): While we assumed $Var(\hat{f}(\boldsymbol{x})^{t}_{j})=s$ for simplicity, in reality, edge models that perform poorly on the validation set or are highly uncertain might themselves exhibit higher intrinsic variance or be more influenced by noise in their data subset.
+By down-weighting these models (assigning smaller $\alpha_{j}$), DUA-D2C effectively reduces their contribution to the overall variance of the central model.
+If an edge model $M_{j}$ is deemed unreliable (low $\alpha_{j}$), its $s\cdot\alpha_{j}^{2}$ term becomes small.
+  • Reducing Impact of Correlated ”Bad” Models: If several edge models are similarly poor (e.g., all overfit to noise in a similar way, leading to correlated errors, high $c$), but are identified as such by the validation process, their $\alpha_{i}\alpha_{l}$ terms will be small, thus mitigating the negative impact of correlated, poorly generalizing models.
+  • Optimal Weighting for Variance Reduction: From a portfolio theory perspective in statistics, for a given set of expected returns (analogous to model quality) and a covariance matrix, there exists an optimal set of weights that minimizes variance for a target return.
+While DUA-D2C’s heuristic weighting (based on validation accuracy and uncertainty) is not explicitly solving such an optimization problem in each step, it intuitively steers the aggregation towards a combination that should exhibit lower variance than a simple average, especially if the quality of edge models varies significantly.
+If all $\alpha_{j}=1/k$, Eq.
+(16) reduces to Eq.
+(13).
+However, if DUA-D2C can assign larger weights to genuinely better, less noisy models, the term $\sum\alpha_{j}^{2}$ can be smaller than $1/k$ (since $\sum\alpha_{j}=1$ and $\alpha_{j}\geq 0$, $\sum\alpha_{j}^{2}\leq(\sum\alpha_{j})^{2}=1$; for uniform weights, $\sum(1/k)^{2}=k/k^{2}=1/k$).
+This suggests a potential for further variance reduction compared to simple D2C if the weighting scheme is effective.
+Essentially, DUA-D2C attempts to perform a ”smarter” averaging.
+Instead of treating all edge models equally, it gives preference to those that demonstrate better generalization capabilities on data not seen during their local training.
+This targeted aggregation is hypothesized to lead to a central model that not only benefits from the diversity of the D2C approach but also more robustly distills the ”signal” from the ”noise” by selectively emphasizing more reliable constituent models.
+The goal is to achieve a central model $\hat{f}(\boldsymbol{x})^{t+1}$ that is closer to the true underlying function $f_{0}$ by minimizing not just variance but also potentially bias, if ”better” edge models are indeed less biased.
+In summary, the D2C framework, by averaging parameters from models trained on data shards, offers a mechanism to reduce the variance of the final aggregated model, particularly when edge models exhibit diversity due to the distribution of noise and outliers.
+The DUA-D2C extension further refines this by employing a dynamic, uncertainty-aware weighting scheme.
+This intelligent aggregation aims to preferentially incorporate knowledge from more reliable and generalizable edge models, potentially leading to even greater variance reduction and a central model that is more robust to overfitting.
+While this analysis relies on simplified assumptions for complex neural networks, it provides a foundational mathematical intuition for why these ”divide and conquer” strategies, especially with adaptive aggregation, are promising for improving model generalization.
+3. Proposed Methodology
+The proposed Dynamic Uncertainty-Aware Divide2Conquer (DUA-D2C) framework enhances traditional distributed training paradigms by introducing an intelligent aggregation mechanism.
+The methodology can be broken down into distinct phases: data partitioning, edge model training, edge model evaluation and weighting, and finally, dynamic global aggregation.
+3.1. Creating Training Subsets
+The initial phase involves preparing the data for distributed processing.
+The entire training dataset $D$ is first randomly shuffled to ensure an unbiased distribution of samples.
+Subsequently, $D$ is partitioned into $N$ distinct, non-overlapping subsets or shards: $\{D_{1},D_{2},\dots,D_{N}\}$.
+A critical consideration during this partitioning is the maintenance of class balance; efforts are made to ensure that the class proportions within each subset $D_{i}$ are representative of the overall dataset $D$.
+Each subset $D_{i}$ is then assigned to a dedicated edge model, $M_{i}$.
+All $N$ edge models, along with a central model $M_{c}$, share an identical neural network architecture and are initialized with the same trainable parameters to facilitate coherent parameter aggregation.
+3.2. Local Training of Edge Models
+Once the data subsets are prepared and edge models initialized (typically with parameters from the central model of the previous global epoch, or randomly at the start), the local training phase commences.
+This phase constitutes the inner loop of the DUA-D2C process.
+Each edge model $M_{i}$ is trained exclusively on its assigned data subset $D_{i}$ for a predetermined number of local epochs ($E$).
+Standard training procedures, such as mini-batch stochastic gradient descent with a given batch size $B$ and learning rate $lr$, are employed during these local epochs.
+Upon completion of $E$ local epochs, each edge model $M_{i}$ possesses an updated set of parameters, $\theta^{\prime}_{i}$.
+3.3. Edge Model Evaluation and Dynamic Weighting
+This phase is central to the DUA-D2C approach and distinguishes it from simpler aggregation schemes.
+After local training, each edge model $M_{i}$ (with its updated parameters $\theta^{\prime}_{i}$) is evaluated on a common, independent validation dataset, $D_{val}$.
+This evaluation serves two purposes:
+  1. (1)
+    To assess the generalization performance of $M_{i}$ via its validation accuracy, $a_{i}$.
+  2. (2)
+    To quantify the predictive uncertainty of $M_{i}$, typically measured by its average prediction entropy, $\epsilon_{i}$, over all samples in $D_{val}$.
+The raw inverse entropy, $inv\_e_{i}=1/(\epsilon_{i}+\delta_{e})$ (where $\delta_{e}$ is a small constant to prevent division by zero), is first computed.
+This value is then capped at a predefined maximum $C_{max}$ to prevent undue influence from overly confident models: $inv\_e_{cap_{i}}=\min(inv\_e_{i},C_{max})$.
+This capped inverse entropy is subsequently Min-Max scaled to a range of $[0,1]$ to represent a normalized confidence score, $u_{i}$:
+(17) $u_{i}=\frac{inv\_e_{cap_{i}}-inv\_e_{min}}{C_{max}-inv\_e_{min}+\epsilon_{s}}$
+where $inv\_e_{min}=1/(\log(K)+\delta_{e})$ is the theoretical minimum inverse entropy (for $K$ classes) and $\epsilon_{s}$ is a small stability constant.
+The value of $u_{i}$ is clamped to ensure it strictly lies within $[0,1]$.
+Finally, a composite raw scaling factor, $s_{i}$, for each edge model $M_{i}$ is calculated as a weighted sum of its validation accuracy and its normalized confidence score:
+(18) $s_{i}=\lambda\cdot a_{i}+(1-\lambda)\cdot u_{i}$
+Here, $\lambda$ is a tradeoff hyperparameter balancing the importance of accuracy versus confidence.
+3.4. Uncertainty-Aware Global Aggregation and Parameter Update
+The global aggregation phase constitutes the outer loop of DUA-D2C.
+The raw scaling factors $\{s_{1},s_{2},\dots,s_{N}\}$ computed in the previous step are first normalized to ensure their sum is unity, yielding the final aggregation weights $\alpha_{i}$ for each edge model $M_{i}$:
+(19) $\alpha_{i}=\frac{s_{i}}{\sum_{j=1}^{N}s_{j}+\epsilon_{s}}$
+The parameters of the central model, $\theta_{c}$, are then updated by taking a weighted average of the parameters $\theta^{\prime}_{i}$ from all $N$ edge models, using these dynamically determined weights $\alpha_{i}$:
+(20) $\theta_{c}\leftarrow\sum_{i=1}^{N}\alpha_{i}\cdot\theta^{\prime}_{i}$
+This weighted aggregation ensures that edge models that demonstrated superior generalization (higher accuracy) and greater reliability (higher confidence) on the validation set contribute more significantly to the updated central model.
+After the central model $M_{c}$ is updated with the new $\theta_{c}$, its parameters are disseminated back to initialize all edge models for the subsequent global epoch.
+This entire cycle of local training, evaluation, dynamic weighting, and global aggregation is repeated for a specified number of global epochs ($E_{global}$).
+The detailed pseudo-code for the complete DUA-D2C method is provided in Algorithm 1.
+Algorithm 1 Dynamic Uncertainty-Aware Divide2Conquer (DUA-D2C)
+1:Input: Training data $D$, Validation data $D_{val}$; Num. subsets NN; Epochs $E,E_{global}$; Batch $B$; Learning rate lrlr; Parameters: $\lambda,C_{max},\delta_{e},K,\epsilon_{s}$
+2:Output: Central model $M_{c}$ with aggregated weights $\theta_{c}$
+3:
+4:function DUA-D2C($D,D_{val},N,E,E_{global},B,lr,\lambda,C_{max},\delta_{e},K,\epsilon_{s}$)
+5:     Shuffle $D$; Divide $D$ into NN subsets: $\{D_{1},\dots,D_{N}\}$ maintaining class distributions
+6:     Initialize central model $M_{c}$ with weights $\theta_{c}$
+7:     Set min. inverse entropy: $inv\_e_{min}=1/(\log(K)+\delta_{e})$
+8:for $g=1$ to $E_{global}$ do
+9:      $W_{central}=0$; $S_{raw}=[\,]$
+10:      Let $\Theta^{\prime}_{edge}=[\,]$ be a list to store edge model weights
+11:  for $i=1$ to NN do
+12:        Initialize edge model $M_{i}$ with $\theta_{i}=\theta_{c}$
+13:   for $e=1$ to $E$ do
+14:          Train $M_{i}$ on $D_{i}$ (batch $B$, learning rate lrlr)
+15:   end for
+16:
+17:        Compute $a_{i}$ (accuracy of $M_{i}$ on $D_{val}$)
+18:        Compute $\epsilon_{i}$ (prediction entropy of $M_{i}$ on $D_{val}$)
+19:        $inv\_e_{i}=1/(\epsilon_{i}+\delta_{e})$
+20:        $inv\_e_{cap_{i}}=\min(inv\_e_{i},C_{max})$
+21:        $u_{i}=(inv\_e_{cap_{i}}-inv\_e_{min})/(C_{max}-inv\_e_{min}+\epsilon_{s})$
+22:        $u_{i}=\max(0,\min(1,u_{i}))$ {Clamp to [0,1][0,1]}
+23:        $s_{i}=\lambda\cdot a_{i}+(1-\lambda)\cdot u_{i}$
+24:        Append $s_{i}$ to $S_{raw}$
+25:        Obtain updated weights $\theta^{\prime}_{i}$ from $M_{i}$; Append $\theta^{\prime}_{i}$ to $\Theta^{\prime}_{edge}$
+26:  end for
+27:
+28:      $Sum\_S_{raw}=\sum_{j=1}^{N}S_{raw}[j]$
+29:      Let $S_{norm}=[\,]$ be a list for normalized scaling factors
+30:  for $i=1$ to NN do
+31:        $S_{norm}[i]=S_{raw}[i]/(Sum\_S_{raw}+\epsilon_{s})$
+32:  end for
+33:
+34:  for $i=1$ to NN do
+35:        $\theta_{scaled_{i}}=S_{norm}[i]\cdot\Theta^{\prime}_{edge}[i]$
+36:        $W_{central}=W_{central}+\theta_{scaled_{i}}$
+37:  end for
+38:
+39:      $\theta_{c}=W_{central}$ {Update central model weights}
+40:      Update $M_{c}$ with $\theta_{c}$
+41:end for
+42:
+43:return central model $M_{c}$
+44:end function
+3.5. Global Hyperparameter Tuning for DUA-D2C
+Tuning the DUA-D2C model involves a structured approach.
+First, the base architecture for the edge models (and thus the central model) and its associated learning hyperparameters (e.g., learning rate, batch size for local training) should be reasonably optimized.
+This can be done by training a single instance of this architecture on a representative portion of the training data or the entire training set initially, using the validation set for guidance.
+Once a suitable base model architecture is established, the DUA-D2C specific global hyperparameters require tuning. These include:
+  • Number of Subsets ($N$): The training data is divided into $N$ shards. The optimal $N$ depends on dataset size, complexity, and class distribution.
+  • Number of Local Epochs ($E$): The number of epochs each edge model trains on its subset before aggregation.
+A small number (e.g., $1\sim 3$) is often preferred, as edge models can quickly overfit their reduced-size training subset.
+  • Tradeoff Parameter ($\lambda$): This parameter balances the contribution of validation accuracy ($a_{i}$) versus scaled confidence ($u_{i}$) in determining the edge model weights.
+Its range is typically $[0,1]$.
+In our experiments, we found the range $0.5\sim 0.8$ to be effective.
+  • Max Inverse Entropy Cap ($C_{max}$): This caps the raw inverse entropy value before scaling, preventing extremely confident (but potentially incorrect) models from dominating.
+This value may be dataset-dependent.
+  • Other parameters like $\delta_{e}$ (entropy smoothing) and $\epsilon_{s}$ (scaling stability) are typically small constants and less frequently tuned, but their values should be noted.
+The tuning process can proceed by first determining a sensible range for $N$ and $E$.
+One might start by setting $E$ to a small value (e.g., 1 or 2) and then varying $N$, observing performance metrics like validation accuracy/loss, F1 score, etc., of the final central model.
+Once a reasonable $N$ is found, $E$ can be tuned.
+Subsequently, $\lambda$ can be optimized by observing the performance metrics too.
+$C_{max}$ can be set to a reasonable value solely by observing the typical range of inverse entropy for a particular dataset.
+The objective is to find the combination that yields the best generalization.
+The most influential hyperparameter among these four is the number of subsets $N$, as we can see from the mathematical intuition.
+4. Experimental Setup
+4.1. Datasets
+We selected a range of benchmark datasets from three different domains to test and validate our proposed method.
+For image classification tasks, we employed three widely used benchmark datasets: CIFAR-10 (krizhevsky2014cifar, 15), MNIST (lecun1998gradient, 16), and Fashion MNIST (xiao2017fashion, 29).
+CIFAR-10 provides a challenging image classification task with 60,000 $32\times 32$ RGB images across 10 categories.
+MNIST offers a simpler task with 70,000 $28\times 28$ grayscale images of handwritten digits.
+Fashion MNIST presents a more complex variation with 70,000 $28\times 28$ grayscale images of fashion items.
+To address class imbalance and the overfitting challenge in facial expression recognition, we utilized the FER-2013 (zahara2020facial, 31) dataset, which comprises over 35,000 grayscale images of seven basic emotions.
 The diverse and noisy nature of this internet-sourced dataset makes it particularly suitable for evaluating generalization performance.
-For audio-based tasks, we created a comprehensive dataset by combining TESS [19], CREMA-D [20], and RAVDESS [21].
+For audio-based tasks, we created a comprehensive dataset by combining TESS (dupuis2010toronto, 23), CREMA-D (cao2014crema, 4), and RAVDESS (livingstone2018ryerson, 18).
 This dataset encompasses 14 emotion classes, offering a diverse range of emotional expressions and recording conditions.
 By merging these datasets, we aim to improve model generalization by mitigating biases and limitations inherent in individual datasets.
-To test our method on text data, we employed the AG News [22] dataset, which consists of approximately 127,600 news articles categorized into four balanced classes: World, Sports, Business, and Science/Technology.
+To test our method on textual data, we employed the AG News (zhang2015character, 33) dataset, which consists of approximately 127,600 news articles categorized into four balanced classes: World, Sports, Business, and Science/Technology.
 This dataset provides a large-scale, real-world benchmark for evaluating text classification models, particularly in the context of Big Data.
-V-B Experimental Details
+4.2. Experimental Details
 1.
@@
 Before constructing the model, processing each training subset into tensors, and batching them was our last step.
-Each of our edge models was comprised of 4 Convolutional (32 to 256 kernels progressively) and MaxPooling layer (2X2) blocks followed by one fully connected hidden layer (128 neurons).
+Each of our edge models was comprised of 4 Convolutional (32 to 256 kernels progressively) and MaxPooling layer (2X2) blocks, followed by one fully connected hidden layer (128 neurons).
 Each block contained two identical convolutional layers and one pooling layer.
@@
 In all cases, we used a 90-10 Training–Validation split, Adam optimizer, and applied dropouts and batch normalization layers.
-This is important to test for D2C’s effectiveness when used on top of other methods that address overfitting.
+This is important to test for DUA-D2C’s effectiveness when used on top of other methods that address overfitting.
 We used the predefined test sets for testing.
-VI Results and Discussions
-We will use loss and accuracy curves, Accuracy, F1 Score, AUC-ROC, and log loss to analyze the results and evaluate our method.
+5. Results and Discussions
+We will use loss and accuracy curves, Accuracy, F1 Score, MCC, Cohen’s Cappa, AUC-ROC, and log loss to analyze the results and evaluate our method.
 We also varied the two new global hyperparameters we mentioned in the previous section: the Number of Subsets of the Training Set and the Number of Epochs before each round of Central Averaging.
 To refer to them concisely, we will use the variables $\boldsymbol{N}$ and $\boldsymbol{E}$ respectively in this section.
-(a) No Subsets (Traditional)
-(b) N = 3 (D2C)
-Figure 4: Decision boundary Visualization with the Traditional approach vs. the Divide2Conquer method.
-Figure 5: Decision Boundary from the edge models trained on each subset after dividing the training data into 3 subsets applying the D2C method.
-VI-A Visualizing Decision Boundary
-At first, we will visualize the change in decision boundaries due to applying the D2C method and analyze the findings.
+As for the Tradeoff Parameter ($\lambda$), we used a value of 0.7 in each case for the sake of consistency, and due to the fact that we found this value to be a good one.
+5.1. Visualizing Decision Boundary
+At first, we will visualize the change in decision boundaries due to applying our proposed method and analyze the findings.
 We used a synthetic binary classification dataset to simulate a simple decision boundary.
@@
 The neural network architecture we used for these experiments comprised three hidden layers having 100 neurons each.
-No regularization or data augmentation was used for decision boundary visualization, to observe the effect of D2C explicitly.
-At first, we adopted the centralized, traditional approach and fed the training dataset to the sole model. After training, we got the decision boundary shown in Figure 4a.
+No regularization or data augmentation was used for decision boundary visualisation, to observe the effect of the Divide2Conquer approach explicitly.
+For decision boundary analysis, we kept things simple and used the basic D2C approach without performance-based aggregation.
+At first, we adopted the centralized, traditional approach and fed the training dataset to the sole model. After training, we got the decision boundary shown in Figure LABEL:fig:_n1-dbl.
+(a) No Subsets (Traditional)
+(b) N = 3 (D2C)
+Figure 3. Decision boundary visualization with the Traditional approach vs. Divide2Conquer (D2C).
+Figure 4. Decision Boundary from the edge models trained on each subset after dividing the training data into 3 subsets using the proposed method.
 As we can see, the decision boundary for the traditional approach is highly complex and non-linear here.
 The model creates intricate regions to distinguish between the two classes.
-The boundary closely follows individual points, resulting in a jagged, convoluted shape.
+The boundary closely follows individual points, resulting in a convoluted, jagged shape.
 The complexity of the decision boundary suggests that the model is overfitting to the training data.
@@
 The highly irregular boundary reflects poor generalization ability.
-Then, to see the effect of D2C on decision boundaries, we divided our training set into multiple subsets and trained them parallelly using identical Edge Models.
+Then, to see the effect of D2C on decision boundaries, we divided our training set into multiple subsets and trained them in parallel using identical Edge Models.
 Figure 4 shows the decision boundaries we got from each edge model.
@@
 Hence, after averaging the weights, the impact of the individual outliers should be diluted.
-To verify, the trained weights from each subset were averaged to get the final weights for the Central Model. The resulting decision boundary is shown in Figure 4b.
-We can see that the decision boundary from the central model of D2C is much simpler, dividing the feature space almost diagonally.
+To verify, the trained weights from each subset were averaged to get the final weights for the Central Model. The resulting decision boundary is shown in Figure LABEL:fig:_n3-db.
+We can see that the decision boundary from the central model of our method is much simpler, dividing the feature space almost diagonally.
 This suggests a less flexible, more generalized model due to averaging the weights from multiple subsets.
@@
 Further proof of this generalization is seen in the training and test accuracies for the traditional approach and D2C using different numbers of subsets.
-We summarize these accuracies in Table I.
-TABLE I: Performance of the traditional approach vs different values of $N$ Using the D2C method.
+We summarize these accuracies in Table 1.
+Table 1. Performance of the traditional approach vs different values of NN Using the proposed method.
 Model Settings Training Accuracy Test Accuracy
@@
 However, several other techniques address overfitting.
-So, whether using D2C on top of other regularization and augmentation techniques (such as dropouts) improves the performance of a model is the key question in application level.
+So, whether using our proposed method on top of other regularization and augmentation techniques (such as dropouts) improves the performance of a model is the key question at the application level.
 In the following subsections, we will investigate this using the architectures (incorporating regularization methods like dropouts in the edge models) we described in the previous section on several datasets from different domains.
-VI-B Analyzing Loss & Accuracy Curves
+For all these experiments, we used the refined DUA-D2C approach.
+5.2. Analyzing Loss & Accuracy Curves
 We will analyze the results obtained from our experiments from the 3 different domains.
@@
 (6), The validation/test loss truly represents the true error of the model.
-So we will monitor the validation loss curves from the different experiments we performed to compare and analyze if D2C improves the generalizing performance, in the form of minimizing the validation loss as the training progresses.
+So we will monitor the validation loss curves from the different experiments we performed to compare and analyze if DUA-D2C improves the generalizing performance, in the form of minimizing the validation loss as the training progresses.
 Image Classification: Let us first take a look at the loss curves generated using the CIFAR-10, Fashion MNIST, MNIST, and FER2013 datasets.
-For now, we will keep the number of epochs before each round of central averaging, $\boldsymbol{E}$, equal to $1$.
+For now, we will keep the number of epochs before each round of central aggregation, $\boldsymbol{E}$, equal to $1$.
 It will ensure a proper comparison between the traditional method and our proposed method while varying the number of subsets, $\boldsymbol{N}$.
 In our experiments, how much we varied $N$ depended on the performance (Loss curve, Accuracy) trend as we increased $N$, the number of subsets.
-In the case of all the datasets, we stopped varying $N$ whenever we observed a definitive performance drop with increasing N.
+In the case of all the datasets, we stopped varying $N$ whenever we observed a definitive and consistent performance drop with increasing $\boldsymbol{N}$.
 For example, in the case of CIFAR-10, we varied $N$ in the range of $1$ to $9$.
-Figure 6: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on CIFAR-10.
-Figure 7: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on Fashion MNIST.
-Figure 8: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on MNIST.
-Figure 9: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on FER2013.
-As we can see from Figure 6 and Figure 9, the validation loss keeps decreasing with each epoch for longer when using our proposed D2C method on these datasets but starts to increase after very few epochs again in the case of traditional NN implementations using the entire training set.
-In Figure 7, we can see that D2C with 9 subsets almost entirely eliminates the trend of an increasing loss after a few epochs.
-In machine learning, overfitting is marked by a model’s increasing validation loss after a certain number of training epochs, indicating that the model has started to capture noise and specific variations in the training set that do not generalize to new data.
-A model that reaches this overfitting point later, meaning it maintains a lower validation loss over a longer period, suggests it is better regularized or has a complexity that is well-suited to the underlying patterns in the data.
-A model with a delayed overfitting point continues to find generalizable patterns within the data through many epochs, highlighting that it remains in a generalization phase for longer, rather than immediately entering an overfitting phase.
-We can also see that, in all these cases, dividing the training data into more and more subsets delays and slows down the abrupt increase in validation loss as the training progresses.
-For example, in the case of CIFAR-10, the loss curve generated using the traditional approach without dividing the training data (no subsets) reaches its minimum, $0.536$, at around $16$ epochs.
-Then it keeps increasing and reaches $0.77$ by the end of 100 epochs.
-As we set the number of epochs before each round of central averaging, $E=1$, each central epoch is equivalent to one epoch in the traditional approach.
-For the number of subsets, $N=3$, the loss curve reaches its minimum, $0.462$, at around $21$ epochs and then it keeps increasing and reaches0.640.640.640.64 by the end of 100 epochs.
-When we further divided the training data into more subsets, setting $N=9$, the loss curve reaches its minimum,$0.47$, at around 46 epochs and then it keeps increasing and reaches $0.585$ by the end of 100 epochs.
-So, clearly, the convergence becomes slower, and the abrupt increase of validation loss is minimized as we divide the training set into more and more subsets, which indicates reduced overfitting.
-Also, D2C manages to maintain a lower validation loss in general in all cases, including MNIST as we can see from Figure 8.
-The minimum value of loss that each of the cases discussed above also tells a similar story.
-The more subsets we divide the training set into, the less the validation loss we can observe from the curves.
-When a model maintains a lower validation loss during training, it indicates that the model is able to generalize well to unseen data.
-Theoretically, overfitting occurs when a model becomes overly complex relative to the data it is learning, allowing it to capture not only meaningful patterns but also specific noise or minor irregularities unique to the training set.
-This focus on noise causes the model to perform poorly on new data, which is reflected in a higher validation loss or an accelerated rise in validation loss.
-In contrast, a model that keeps validation loss lower while training likely has a structure that is well-suited to the data’s true underlying patterns without becoming excessively sensitive to its unique characteristics.
-By maintaining a lower validation loss, the model demonstrates an ability to generalize—learning patterns that extend beyond the training set and apply well to validation data.
-This suggests that the model has achieved a balance, where it is complex enough to capture essential relationships within the data but not so complex that it becomes reliant on details that do not generalize.
-Consequently, a model that sustains a lower validation loss is theoretically overfitting less, as it shows a stronger alignment with the data’s general structure rather than its idiosyncrasies.
-So, we can conclude that our proposed method manages to bring down the validation loss during training compared to the traditional method, which is a definitive sign of better generalization and reduced overfitting.
-We can further conclude that this reduction in validation loss becomes more apparent as we increase $N$, the number of subsets.
-The most clear illustration of this phenomenon is seen in the case of FER2013 (Figure 9).
-The Facial Expression Recognition task using the FER2013 dataset is prone to overfitting to a great extent.
-Dividing the training set into an increasing number of subsets evidently minimizes that overfitting.
-So, dividing the training set into more and more subsets helps reduce overfitting, an observation that is consistent with what we saw in Eq.
+Figure 5. Validation loss curves for the traditional & the DUA-D2C method for different numbers of subsets, N, on CIFAR-10.
+Figure 6. Validation loss curves for the traditional & the DUA-D2C method for different numbers of subsets, N, on Fashion MNIST.
+Figure 7. Validation loss curves for the traditional & the DUA-D2C method for different numbers of subsets, N, on MNIST.
+Figure 8. Validation loss curves for the traditional & the DUA-D2C method for different numbers of subsets, N, on FER2013.
+As we can see in Figures 5, 6, and 8, a key indicator of overfitting, an increasing validation loss after initial training, is notably delayed and mitigated when using our proposed DUA-D2C method compared to traditional implementations.
+In machine learning, overfitting is often marked by a model’s increasing validation loss after a certain number of training epochs, indicating that the model has started to capture noise and specific variations in the training set that do not generalize.
+A model that reaches this overfitting point later, as DUA-D2C does, suggests it remains in a generalization phase for longer.
+We can also see that, in all these cases, dividing the training data into more subsets with DUA-D2C delays and slows down the abrupt increase in validation loss.
+For example, on CIFAR-10, the traditional approach (no subsets) reached its minimum loss of $0.536$ at around $16$ epochs, then increased to $0.77$ by 100 epochs.
+With DUA-D2C ($E=1$), for $N=3$, the minimum loss was $0.462$ around $21$ epochs, rising to $0.64$.
+When $N=9$, the minimum was $0.47$ around 46 epochs, reaching $0.585$.
+So, clearly, the convergence to the overfitting point becomes slower, and the abrupt increase of validation loss is minimized as we divide the training set into more subsets, which indicates reduced overfitting with DUA-D2C.
+Also, DUA-D2C manages to maintain a lower validation loss in general in all cases, including MNIST, as we can see from Figure 7.
+The minimum value of loss achieved also tells a similar story: we can observe the minimum validation loss for a DUA-D2C variant in each case.
+When a model maintains a lower validation loss, it indicates better generalization, as it effectively captures true underlying patterns without becoming excessively sensitive to training set noise.
+This stronger alignment with the data’s general structure rather than its idiosyncrasies means the model is overfitting less.
+The most striking illustration of this phenomenon is seen with the FER2013 dataset (Figure 8), a task highly prone to overfitting.
+Here, increasing the number of subsets with DUA-D2C clearly diminishes this overfitting tendency.
+This empirical evidence, that employing more subsets helps reduce overfitting, aligns with our theoretical intuition from Eq.
 (13).
-The more the number of subsets ($k$ in that equation), the less the variance, s/k, for the best-case scenario.
-Hence, with the decrease in variance comes the decrease in overfitting.
-Also, as we saw before, the true error, represented by validation loss in this case, is minimized using D2C, which proves our hypothesis and shows that D2C is indeed effective in treating the problem of overfitting.
-Figure 10: Validation accuracy curves for the traditional method & the D2C method for different numbers of subsets, N, on Fashion MNIST.
+As $k$ (number of subsets) increases, the variance of the aggregated model can decrease (approaching $s/k$ in the ideal scenario), directly combating a primary component of overfitting.
+Thus, the observed minimization of validation loss (our proxy for true error) with DUA-D2C, particularly with a larger $N$, validates our hypothesis and confirms its effectiveness in treating the problem of overfitting.
+There is another interesting aspect of the validation loss curve here.
+As we can see from Figure 7, the validation loss curve generated using the traditional approach deviates much more than the ones generated using the DUA-D2C method.
+A curve that does not fluctuate much after converging suggests that the model has reached a stable state.
+This stability is often a sign of robustness, meaning the model is not overly sensitive to slight variations in the data or training process.
+Dividing the dataset into more and more subsets results in smaller and smaller deviations in the validation loss curve.
+This was evident in previous loss curves too, to a lesser extent.
+This proves that dividing the training set into multiple subsets not only improves accuracy and loss but also brings stability and consistency to the loss/accuracy curves.
+Figure 9. Validation accuracy curves for the traditional & the DUA-D2C method for different numbers of subsets, N, on Fashion MNIST.
 However, the loss curve isn’t the only important factor during the training process.
@@
 This shows why continuing training for a significant period might be necessary even if the validation loss starts increasing.
-This also shows why early stopping based on loss curves may not be a good idea in both cases: the traditional approach and the D2C method.
+This also shows why early stopping based on loss curves may not be a good idea in both cases: the traditional approach and the DUA-D2C method.
 This is something we observed in the cases of all these datasets.
@@
 At the very least, dividing the training set into multiple subsets ensures that the maximum accuracy is attained while keeping the log loss relatively in control, which characterizes a more confident and robust model with better generalization ability.
-There is another interesting aspect of the validation loss curve here.
-As we can see from Figure 8, the validation loss curve generated using the traditional approach deviates much more than the ones generated using the D2C method.
-A curve that does not fluctuate much after converging suggests that the model has reached a stable state.
-This stability is often a sign of robustness, meaning the model is not overly sensitive to slight variations in the data or training process.
-Dividing the dataset into more and more subsets results in smaller and smaller deviations in the validation loss curve.
-This was evident in previous loss curves too, to a lesser extent.
-This proves that dividing the training set into multiple subsets not only improves accuracy and loss but also brings stability and consistency to the loss/accuracy curves.
-Let us also take a look at the validation loss curves generated using FER2013 while keeping the number of subsets, $N$, constant and varying the number of local epochs before each round of central averaging, $E$.
+Let us also take a look at the validation loss curves generated using FER2013 while keeping the number of subsets, $N$, constant and varying the number of local epochs before each round of central aggregation, $E$.
 We varied $E$ between $1$ to $3$ and Kept $N$ constant at $2$ and $3$ separately.
-Figure 11: Validation loss curves for different numbers of epochs before each round of central averaging, while keeping the number of subsets, N = 2 & 3.
-We can see from Figure 11 that when we vary the number of epochs before each round of central averaging, $E$, there is no clear indication of any kind of change in the trend or the rate of change in the validation losses as the training progresses.
-This is unlike when we varied the number of subsets, $N$ keeping $E$ constant, where we saw that increasing the number of subsets delayed and minimized the increase in validation loss as the training progressed quite clearly.
+Figure 10. Validation loss curves for different numbers of epochs before each round of central averaging, while keeping the number of subsets, N = 2 & 3.
+We can see from Figure 10 that when we vary the number of epochs before each round of central aggregation, $E$, there is no clear indication of any kind of change in the trend or the rate of change in the validation losses as the training progresses.
+This is unlike when we varied the number of subsets, $N$, keeping $E$ constant, where we saw that increasing the number of subsets delayed and minimized the increase in validation loss as the training progressed quite clearly.
 Also, the convergence got slower.
 But here, for both $N=2$ (shades of blue) and $N=3$ (shades of green), the trends from the validation curves are quite similar for different values of $E$ throughout the training phase.
-In light of the above observations, we can conclude that varying the number of epochs before each round of central averaging, $E$ does not have any obvious implications on log loss as we saw in the case of varying the number of subsets, $N$.
-However, an increased number of local epochs before each round of central averaging means more scope for learning from each of the subsets before averaging, sometimes even overfitting on a particular subset a bit more.
+In light of the above observations, we can conclude that varying the number of epochs before each round of central aggregation, $E$ does not have any obvious implications on log loss as we saw in the case of varying the number of subsets, $N$.
+However, an increased number of local epochs before each round of central aggregation means more scope for learning from each of the subsets before combining and assigning the weights, sometimes even overfitting on a particular subset a bit more.
 We will observe the effect of varying the number of epochs further in the later subsections when we observe the final accuracy and log loss in each case.
@@
 One way of tackling this issue is by combining datasets to increase the diversity in both training and test/validation data.
-We combined the datasets TESS [19], CREMA-D [20], and RAVDESS [21] for this task.
+We combined the datasets TESS (dupuis2010toronto, 23), CREMA-D (cao2014crema, 4), and RAVDESS (livingstone2018ryerson, 18) for this task.
 Let us take a look at the validation loss and accuracy curves generated using this Combined Audio dataset.
-Figure 12: Validation loss curves for the traditional method & the D2C method for different numbers of subsets, N, on the Combined Audio dataset.
+Figure 11. Validation loss curves for the traditional & the DUA-D2C method for different numbers of subsets, N, on the Combined Audio dataset.
 As we can see from Figure 11, there is overfitting to some extent in all cases, which causes the validation loss to creep up after a few epochs but the rate at which this phenomenon takes place is much less when the training set is divided into multiple subsets.
@@
 Even though we converted the 1D time series audio files to 2D Log-Mel spectrograms to form something like an image, the classification based on these Mel-specs differs greatly from traditional image classification.
-So, the overfitting-resistant nature of models using our proposed method being apparent in this case too is a very promising sign.
-However, the small size of this dataset may not be ideal in the context of applying D2C.
+So, the overfitting-resistant nature of models using our proposed method being apparent in this case is a very promising sign.
+However, the small size of this dataset may not be ideal in the context of applying DUA-D2C.
 We will analyze it in detail in the next subsection based on several performance metrics.
-Text Classification: Text classification is an important case study when it comes to testing the Divide2Conquer method.
+Text Classification: Text classification is an important case study when it comes to testing the DUA-D2C method.
 For both image classification and audio classification, we used 2D data and CNN to carry out the tasks.
-But text classification deals with 1D sequences, and as we mentioned in the previous section, we will use bi-directional LSTM in this case.
-We used AG News [22], a benchmark dataset to test our method primarily in the NLP domain.
+But text classification deals with 1D sequences, and as we mentioned in the previous section, we used bi-directional LSTM in this case.
+We used AG News (zhang2015character, 33), a benchmark dataset to test our method primarily in the NLP domain.
 Let us observe the validation loss and accuracy curves generated using the AG News dataset first.
-Figure 13: Validation loss curves for the traditional & the D2C method for different numbers of subsets on AG News.
-Figure 14: Validation accuracy curves for the traditional & the D2C method for different numbers of subsets on AG News.
+Figure 12. Validation loss curves for the traditional & the DUA-D2C method for different numbers of subsets on AG News.
+Figure 13. Validation accuracy curves for the traditional & the DUA-D2C method for different numbers of subsets on AG News.
 The loss and accuracy curves generated for AG News depict by far the clearest picture in terms of the improvements brought by the proposed method over the traditional monolithic approach.
@@
 More importantly, with no subsets, the validation loss shoots up sharply after around 8 epochs, which indicates overfitting and impacts the accuracy too as the training progresses as we can see from Figure 13.
-However, when we use D2C, the sharp increase in the validation loss is mitigated drastically.
+However, when we use DUA-D2C, the sharp increase in the validation loss is mitigated drastically.
 Even with only two subsets, the improvement is quite significant.
 When the number of subsets is increased further to more than two, the loss curve becomes quite stable, almost completely eliminating the increasing trend and showing consistent improvement as the training progresses.
-The accuracy curve (Figure 14) also shows clear indications that D2C improves over the traditional approach.
-With the traditional approach, the accuracy too starts deteriorating after around the 14th epoch, indicating severe overfitting, an issue that no longer exists when we use D2C for any value of $N$ or $E$.
-The experiments on all these datasets from 3 different datasets prove that using the D2C method can indeed address overfitting and control the abrupt increase of true error in the form of maintaining a lower validation loss during training.
-This also shows that D2C works well in different domains and a diverse set of tasks.
-VI-C Accuracy, F1 Score, AUC-ROC, and Log Loss
-We evaluated our proposed D2C on test sets using Accuracy, F1 Score, Log Loss, and AUC-ROC to capture a comprehensive view of model performance in terms of correctness, class balance, prediction confidence, and discrimination ability.
+The accuracy curve (Figure 13) also shows clear indications that DUA-D2C improves over the traditional approach.
+With the traditional approach, the accuracy too starts deteriorating after around the 14th epoch, indicating severe overfitting, an issue that no longer exists when we use DUA-D2C for any value of $N$ or $E$.
+The experiments on all these datasets from 3 different datasets prove that using the DUA-D2C method can indeed address overfitting and control the abrupt increase of true error in the form of maintaining a lower validation loss during training.
+This also shows that DUA-D2C works well in different domains and a diverse set of tasks.
+5.3. Accuracy, F1 Score, AUC-ROC, Cohen’s Cappa, Log Loss and MCC
+We evaluated our proposed DUA-D2C on test sets using Accuracy, F1 Score, AUC-ROC, Cohen’s Cappa, Log Loss, and Matthew’s Correlation Coefficient (MCC) to capture a comprehensive view of model performance in terms of correctness, class balance, prediction confidence, and discrimination ability.
 Reporting these final metrics on test sets was essential to confirm that improvements generalized well beyond the training and validation data, verifying true effectiveness in unseen scenarios.
@@
 We used the provided test sets with each of the datasets for proper comparisons.
-In the tables, the top row represents the traditional centralized approach and the rest of the rows represent different variations of the D2C approach.
+In the tables, the top row represents the traditional centralized approach, and the rest of the rows represent different variations of the DUA-D2C approach.
 The bold entities represent the best performance in the tables.
-Image Classification: Let’s take a look at the performance metrics from the benchmark image classification datasets we primarily used using different settings (the traditional approach and the D2C method with different values of $N$ and $E$).
-TABLE II: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on CIFAR-10.
-Model Settings Accuracy F1-Score Log Loss AUC-ROC
-No Subsets 0.8596 0.8591 0.7346 0.98697
-N=2, E=1 0.8611 0.8611 0.7042 0.98693
-N=3, E=1 0.8613 0.8609 0.6682 0.98704
-N=4, E=1 0.8586 0.8583 0.6516 0.98690
-N=5, E=1 0.8552 0.8551 0.6389 0.98628
-N=7, E=1 0.8539 0.8535 0.6198 0.98640
-N=9, E=1 0.8531 0.8525 0.6068 0.98605
-TABLE III: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on Fashion MNIST.
-Model Settings Accuracy F1-Score Log Loss AUC-ROC
-No Subsets 0.9314 0.9315 0.3467 0.99559
-N=2, E=1 0.9323 0.9319 0.3460 0.99549
-N=3, E=1 0.9330 0.9329 0.3179 0.99545
-N=3, E=2 0.9319 0.9316 0.3303 0.99552
-N=3, E=3 0.9337 0.9337 0.3225 0.99590
-N=5, E=1 0.9322 0.9321 0.2692 0.99589
-N=7, E=1 0.9338 0.9334 0.2467 0.99584
-N=7, E=2 0.9305 0.9304 0.2575 0.99600
-N=9, E=1 0.9323 0.9323 0.2341 0.99600
-TABLE IV: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on MNIST.
-Model Settings Accuracy F1-Score Log Loss AUC-ROC
-No Subsets 0.9952 0.9952 0.0258 0.99995
-N=2, E=1 0.9957 0.9956 0.0271 0.99995
-N=3, E=1 0.9958 0.9958 0.0240 0.99998
-N=3, E=2 0.9951 0.9950 0.0250 0.99996
-N=5, E=1 0.9957 0.9956 0.0202 0.99998
-N=7, E=1 0.9951 0.9950 0.0215 0.99997
-N=9, E=1 0.9957 0.9957 0.0179 0.99998
-N=11, E=1 0.9959 0.9959 0.0161 0.99999
-N=11, E=2 0.9958 0.9958 0.0170 0.99998
-As we can see from Table II, Table III, and Table IV, the best accuracy, MCC and AUC-ROC were achieved using the D2C method for all three datasets.
+Image Classification: Let’s take a look at the performance metrics from the benchmark image classification datasets we primarily used using different settings (the traditional approach and the DUA-D2C method with different values of $N$ and $E$).
+Table 2. Performance Metrics with Traditional Approach vs. Different Settings Using the DUA-D2C Method on CIFAR-10.
+Model Settings Accuracy F1-Score ROC-AUC Cohen Kappa Log Loss Matthews CC
+No Subsets 0.8604 0.8601 0.98731 0.8458 0.7198 0.8458
+N=2, E=1 0.8626 0.8628 0.98704 0.8476 0.7002 0.8476
+N=3, E=1 0.8652 0.8648 0.98761 0.8501 0.6491 0.8501
+N=3, E=2 0.8630 0.8624 0.98719 0.8478 0.6215 0.8478
+N=4, E=1 0.8595 0.8591 0.98710 0.8439 0.6416 0.8439
+N=5, E=1 0.8614 0.8613 0.98676 0.8461 0.6347 0.8461
+N=7, E=1 0.8556 0.8553 0.98658 0.8393 0.6126 0.8398
+N=9, E=1 0.8548 0.8542 0.98664 0.8385 0.5876 0.8390
+Table 3. Performance Metrics with Traditional Approach vs. Different Settings Using the DUA-D2C Method on Fashion MNIST.
+Model Settings Accuracy F1-Score ROC-AUC Cohen Kappa Log Loss Matthews CC
+No Subsets 0.9312 0.9313 0.99546 0.9237 0.3492 0.9238
+N=3, E=1 0.9338 0.9336 0.99570 0.9264 0.2865 0.9265
+N=3, E=2 0.9344 0.9340 0.99571 0.9271 0.3006 0.9272
+N=3, E=3 0.9332 0.9329 0.99580 0.9258 0.3008 0.9258
+N=5, E=1 0.9327 0.9323 0.99581 0.9252 0.2618 0.9253
+N=7, E=1 0.9321 0.9321 0.99591 0.9246 0.2383 0.9246
+N=7, E=2 0.9322 0.9319 0.99600 0.9247 0.2377 0.9247
+N=9, E=1 0.9300 0.9299 0.99590 0.9222 0.2314 0.9222
+Table 4. Performance Metrics with Traditional Approach vs. Different Settings Using the DUA-D2C Method on MNIST.
+Model Settings Accuracy F1-Score ROC-AUC Cohen Kappa Log Loss Matthews CC
+No Subsets 0.9952 0.9952 0.99995 0.9947 0.0258 0.9947
+N=2, E=1 0.9954 0.9954 0.99991 0.9949 0.0294 0.9949
+N=3, E=1 0.9955 0.9955 0.99995 0.9950 0.0225 0.9950
+N=3, E=2 0.9953 0.9953 0.99995 0.9948 0.0259 0.9948
+N=5, E=1 0.9957 0.9957 0.99998 0.9952 0.0188 0.9952
+N=7, E=1 0.9960 0.9959 0.99998 0.9956 0.0183 0.9956
+N=9, E=1 0.9963 0.9963 0.99999 0.9959 0.0148 0.9959
+N=11, E=1 0.9953 0.9953 0.99998 0.9948 0.0158 0.9948
+N=11, E=2 0.9959 0.9959 0.99998 0.9954 0.0156 0.9954
+As we can see from Table 2, Table 3, and Table 4, the best accuracy, F1-score, MCC, Cohen’s Cappa, log loss, and AUC-ROC were achieved using the DUA-D2C method for all three datasets.
 For CIFAR-10, the number of subsets, $N$, was $3$ for the best model in terms of accuracy.
-The number of epochs before each round of central averaging, $E$, was just $1$.
-This approach showed an improvement with an accuracy of $86.13\%$ over the traditional approach with no subsets, which yielded an accuracy of $85.96\%$.
-Initially, the accuracy improved as we increased the number of subsets, but suffered a consistent drop after it was further increased past $3$.
+The number of epochs before each round of central aggregation, $E$, was just $1$.
+This approach showed an improvement with an accuracy of $86.52\%$ over the traditional approach with no subsets, which yielded an accuracy of $86.04\%$.
+Initially, the accuracy improved as we increased the number of subsets, but suffered a decreasing trend after it was further increased past $3$.
 As we discussed, dividing the training set into too many subsets may impact training negatively due to a lack of representative samples.
 The experimental outcomes show further evidence of that in this case.
-For Fashion MNIST too, the best accuracy, MCC, and AUC-ROC were achieved using our proposed Divide2Conquer method.
-The number of subsets, $N$, was $7$ in this case for the best model in terms of accuracy.
-The number of epochs before each round of central averaging, $E$, was once again just $1$.
-This approach again showed an improvement with an accuracy of $93.38\%$ over the traditional approach with no subsets, which yielded an accuracy of $93.14\%$.
-An increasing trend is visible in accuracy as we increase the number of subsets till it reaches the best conditions with $N=7$.
-In the case of MNIST too, the best accuracy, MCC, and AUC-ROC were achieved using our proposed method.
-The number of subsets, $N$, was $11$ in this case for the best model.
-The number of epochs before each round of central averaging, $E$, was $1$.
-The improvement of $0.07$ in error rate over the traditional approach is significant for the MNIST dataset.
+For Fashion MNIST too, the best metrics were achieved using our proposed DUA-D2C method.
+The number of subsets, $N$, was once again $3$ in this case for the best model in terms of accuracy.
+The number of epochs before each round of central aggregation, $E$, was $2$ in this case.
+This approach again showed an improvement with an accuracy of $93.44\%$ over the traditional approach with no subsets, which yielded an accuracy of $93.12\%$.
+An increasing trend is visible in ROC-AUC as we increase the number of subsets till it reaches the best conditions with $N=7$.
+In the case of MNIST too, the best metrics were all achieved using our proposed method.
+The number of subsets, $N$, was $9$ in this case for the best model.
+The number of epochs before each round of central aggregation, $E$, was $1$.
+The improvement of $0.11$ in error rate over the traditional approach is very significant for the MNIST dataset.
 In this case, the least amount of log loss is also achieved with the best model in terms of accuracy.
@@
 The decrease in log loss is a result of that phenomenon.
-Even though the best model isn’t usually selected based on log loss values, in the case of similar accuracy (for example, the accuracies for $N=3,E=3$ and $N=7,E=1$ in Fashion MNIST dataset), the best model can be selected based on log loss.
+Even though the best model isn’t usually selected based on log loss values, in the case of similar accuracy, the best model can be selected based on log loss.
 Log Loss evaluates how well the model predicts probabilities, penalizing confident but incorrect predictions, which is critical for understanding model reliability.
@@
 The comparison of subset numbers in the best-performing models across CIFAR-10, Fashion MNIST, and MNIST provides critical insights.
-For CIFAR-10, the optimal model utilized only $3$ training subsets, while the best models for Fashion MNIST and MNIST required $7$ and $11$ subsets, respectively.
-Notably, increasing the subset count beyond $3$ led to declining accuracy for CIFAR-10—a trend not observed with the other datasets.
+For CIFAR-10, the optimal model utilized only $3$ training subsets, while the best models for Fashion MNIST and MNIST required 33 (77 in terms of ROC-AUC) and $9$ subsets, respectively.
+Notably, increasing the subset count beyond $3$ led to significantly declining accuracy for CIFAR-10, a trend not observed with the other datasets.
 This discrepancy is likely due to the distinct characteristics of each dataset.
@@
 This complexity results in the optimal Fashion MNIST model requiring fewer subsets than MNIST, as each subset’s sample size needs to be sufficient for adequate training.
-This analysis suggests a key principle for applying the D2C method effectively: to divide training data into more subsets, datasets must be large enough to allow each subset ample training data.
+This analysis reiterates a key principle for applying the DUA-D2C method effectively: to divide training data into more subsets, datasets must be large enough to allow each subset enough training data.
 Larger datasets benefit more from divided training, requiring more subsets to achieve optimal performance.
-TABLE V: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on FER2013.
-Model Settings Accuracy F1-Score Log Loss AUC-ROC
-No Subsets 0.6498 0.6420 2.0112 0.89934
-N=2, E=1 0.6511 0.6394 1.9101 0.89875
-N=2, E=2 0.6567 0.6444 1.8450 0.9032
-N=2, E=3 0.6490 0.6389 1.8718 0.8975
-N=3, E=1 0.6491 0.6356 1.6540 0.90104
-N=3, E=2 0.6538 0.6465 1.6990 0.89870
-N=3, E=3 0.6539 0.6428 1.7315 0.9002
-N=5, E=1 0.6500 0.6368 1.5157 0.90153
-N=7, E=1 0.6521 0.6432 1.4821 0.90148
+Table 5. Performance Metrics with Traditional Approach vs. Different Settings Using the DUA-D2C Method on FER2013.
+Model Settings Accuracy F1-Score ROC-AUC Cohen Kappa Log Loss Matthews CC
+No Subsets 0.6498 0.6420 0.89934 0.5756 2.0112 0.5759
+N=2, E=1 0.6512 0.6457 0.90262 0.5781 1.6723 0.5783
+N=2, E=2 0.6510 0.6381 0.90360 0.5776 1.5595 0.5777
+N=2, E=3 0.6480 0.6338 0.89610 0.5740 1.6505 0.5750
+N=3, E=1 0.6471 0.6365 0.89910 0.5735 1.6408 0.5739
+N=3, E=2 0.6482 0.6346 0.90088 0.5743 1.6311 0.5749
+N=3, E=3 0.6420 0.6232 0.89670 0.5671 1.4744 0.5682
+N=5, E=1 0.6537 0.6370 0.90528 0.5816 1.1934 0.5818
+N=5, E=2 0.6482 0.6308 0.90020 0.5741 1.4750 0.5747
+N=7, E=1 0.6474 0.6339 0.90244 0.5732 1.3299 0.5743
 FER2013, a challenging dataset for facial expression recognition, represents a task that is particularly prone to overfitting.
-As seen in Table V, our proposed D2C method outperformed the traditional approach and achieved the highest accuracy, MCC, and AUC-ROC, with the best-performing model using only 2 subsets ($N=2$) and 2 training epochs before each central averaging step ($E=2$).
-D2C yielded a notable accuracy improvement, reaching 65.67% compared to 64.98% in the traditional model without subsets.
-Unlike previous datasets, where a single epoch ($E=1$) was optimal, FER2013 benefited from 2 epochs, likely due to the dataset’s complexity, requiring additional training time to capture meaningful features before model averaging.
+As seen in Table 5, our proposed method outperformed the traditional approach and achieved the best accuracy, Cohen’s Cappa, MCC, log loss, and AUC-ROC, with the best-performing model using 5 subsets ($N=5$) and 1 training epoch before each central aggregation step ($E=1$).
+DUA-D2C yielded a notable accuracy improvement, reaching 65.37% compared to 64.98% of the traditional model without subsets.
+However, D2C performed even better with up to 65.67% of accuracy using the best settings.
 A decreasing trend in log loss was observed with more subsets, though no consistent pattern emerged with varying epochs ($E$).
-Additionally, due to FER2013’s class imbalance, accuracy alone may not reliably indicate the best model, so, the F1-score, a metric considering class-wise precision, provides a valuable perspective.
-Interestingly, the optimal F1-score model (3 subsets, $E=2$) differs from the accuracy-optimized model, with D2C achieving an F1 score of 0.6465 versus 0.642 in the baseline model.
-This F1-score improvement underscores D2C’s effectiveness in addressing overfitting, particularly in class-imbalanced, complex tasks.
+FER2013 is a particularly important test case for us as this dataset represents a comparatively imbalanced class distribution.
+Due to FER2013’s class imbalance, accuracy alone may not reliably indicate the best model, so, the F1-score, a metric considering class-wise precision, provides a valuable perspective.
+Interestingly, the optimal F1-score model (2 subsets, $E=1$) differs from the accuracy-optimized model, with D2C achieving an F1 score of 0.6457 versus 0.642 in the baseline model.
+This F1-score improvement underscores DUA-D2C’s effectiveness in addressing overfitting, particularly in class-imbalanced, complex tasks.
 Audio Classification: Now let us take a look at the performance metrics from the Combined Audio dataset.
-TABLE VI: Performance Metrics with Traditional Approach vs. Different Settings Using D2C on the Combined Audio Dataset.
-Model Settings Accuracy F1-Score Log Loss AUC-ROC
-No Subset 0.6953 0.6870 1.3419 0.96570
-N=2, E=1 0.6942 0.6879 1.0811 0.96553
-N=2, E=2 0.6908 0.6819 1.1636 0.9632
-N=3, E=1 0.6920 0.6909 1.2032 0.96416
-N=3, E=2 0.6874 0.6830 1.0033 0.96438
-N=4, E=1 0.6921 0.6804 1.2840 0.96214
-We can see from Table VI that in the Combined Audio dataset, the highest accuracy was achieved using the traditional training approach rather than our proposed D2C method.
-The traditional model attained a 69.53% accuracy, slightly outperforming the D2C approach, which reached 69.42% with $2$ subsets.
-Accuracy declined further as the number of subsets increased, likely due to the smaller dataset size (11,632 samples), limiting effective data availability per subset.
+This is another important case as combining datasets like TESS, CREMA-D, and RAVDESS is likely to lead to Non-IID data, and this one is another imbalanced and smaller dataset.
+Table 6. Performance Metrics with Traditional Approach vs. Different Settings Using DUA-D2C on the Combined Audio Dataset.
+Model Settings Accuracy F1-Score ROC-AUC Cohen Kappa Log Loss Matthews CC
+No Subsets 0.6953 0.6870 0.96570 0.6684 1.3419 0.6692
+N=2, E=1 0.6869 0.6787 0.96226 0.6592 1.7225 0.6601
+N=2, E=2 0.7010 0.6973 0.96360 0.6747 1.5021 0.6751
+N=3, E=1 0.6946 0.6849 0.96284 0.6678 1.4576 0.6681
+N=3, E=2 0.6775 0.6695 0.96560 0.6492 1.0571 0.6504
+N=4, E=1 0.6930 0.6805 0.96423 0.6659 1.3829 0.6662
+N=5, E=1 0.6938 0.6862 0.96523 0.6668 1.2160 0.6673
+We can see from Table 6 that in the Combined Audio dataset, the best metrics were achieved using our proposed DUA-D2C method once again.
+The traditional model attained a 69.53% accuracy, while the DUA-D2C approach, reached 70.10% with $2$ subsets and $E=2$.
+Accuracy declined as the number of subsets increased, likely due to the smaller dataset size (11,632 samples), limiting effective data availability per subset.
 As a result, dividing the dataset into smaller portions likely impacted performance by reducing the sample count for each class, a significant challenge given the dataset’s complexity and diversity, as it combines samples from multiple sources.
-However, despite marginally lower accuracy, the D2C method showed reduced log loss, indicating a trade-off between accuracy and confidence in predictions.
-In cases of comparable accuracy, lower log loss can be preferable as it indicates more calibrated predictions.
-Moreover, given the dataset’s class imbalance, the F1-score is a more reliable metric here.
-The D2C approach yielded a slight improvement in the F1-score (0.6909 vs.
+However, despite declining accuracy, the DUA-D2C method showed reduced log loss as $N$ increased to 3, indicating a trade-off between accuracy and confidence in predictions.
+Given the dataset’s class imbalance, the F1-score is a more reliable metric here.
+The DUA-D2C approach yielded a significant improvement in the F1-score (0.6973 vs.
 0.687), indicating better generalization across classes.
-This suggests that D2C supports robust classification with fewer, more distinctive errors, especially in class-imbalanced, complex tasks like this one.
-In summary, while D2C did not maximize accuracy, it achieved comparable F1 performance with lower log loss, resulting in a model that generalizes well and offers more confident predictions in challenging, class-imbalanced datasets.
+This suggests that DUA-D2C supports robust classification with fewer, more distinctive errors, especially in class-imbalanced, complex tasks like this one.
+However, the traditional approach outperformed both DUA-D2C and D2C in terms of ROC-AUC for this dataset.
+In summary, while DUA-D2C did not maximize ROC-AUC, it achieved better metrics overall, resulting in a model that generalizes well and offers more confident predictions in challenging, class-imbalanced datasets.
 However, dividing the dataset into too many subsets does worsen the performance of the models in this case.
-This analysis leads us to the conclusion that to apply the Divide2Conquer method effectively, the dataset needs to be sufficiently large so that each of the subsets can have enough training data.
-It is however noteworthy that even with a comparatively smaller dataset, our proposed method still shows some performance improvement in some metrics.
+This analysis leads us to the conclusion that to apply the DUA-D2C method effectively, the subsets need to be sufficiently large so that each of them can have enough training data.
+It is however noteworthy that even with a comparatively smaller dataset, DUA-D2C still shows performance improvement in most metrics.
 Text Classification: Finally, let us take a look at the performance metrics from the AG News dataset.
-TABLE VII: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on AG News.
-Model Settings Accuracy F1-Score Log Loss AUC-ROC
-No Subset 0.9243 0.9243 0.2611 0.98764
-N=2, E=1 0.9291 0.9291 0.2556 0.98873
-N=3, E=1 0.9288 0.9288 0.2520 0.98872
-N=3, E=2 0.9275 0.9275 0.2567 0.98852
-N=5, E=1 0.9293 0.9292 0.2412 0.98881
-N=5, E=2 0.9295 0.9296 0.2451 0.98875
-N=7, E=1 0.9271 0.9271 0.2353 0.98916
-We can observe from Table VII that, once again, the highest accuracy and F1-score were achieved with our proposed D2C method.
-The number of subsets, $N$, was $5$ in this case for the best model.
-The number of epochs before each round of central averaging, $E$, was $2$.
-The D2C method showed a significant improvement with an accuracy of $92.95\%$ over the traditional approach with no subsets, which yielded an accuracy of $92.43\%$.
-Using the D2C method decreased the error by more than $0.5\%$ in this case.
-Also, as we can see from Table VII, the trend of a decreasing log loss as we increase the number of subsets is visible here too.
-The best method in terms of AUC-ROC also happens to be the model with the highest number of subsets (7).
+Table 7. Performance Metrics with Traditional Approach vs. Different Settings Using the DUA-D2C Method on AG News.
+Model Settings Accuracy F1-Score ROC-AUC Cohen Kappa Log Loss Matthews CC
+No Subsets 0.9243 0.9243 0.98764 0.8991 0.2611 0.8992
+N=2, E=1 0.9300 0.9300 0.98856 0.9067 0.2519 0.9067
+N=3, E=1 0.9312 0.9312 0.98888 0.9082 0.2378 0.9084
+N=3, E=2 0.9278 0.9277 0.98879 0.9037 0.2411 0.9037
+N=5, E=1 0.9274 0.9273 0.98905 0.9032 0.2416 0.9032
+N=5, E=2 0.9287 0.9287 0.98830 0.9049 0.2656 0.9049
+N=7, E=1 0.9288 0.9288 0.98836 0.9051 0.2443 0.9052
+We can observe from Table 7 that, once again, the highest accuracy and F1-score were achieved with our proposed DUA-D2C method.
+The number of subsets, $N$, was $3$ in this case for the best model.
+The number of epochs before each round of central averaging, $E$, was $1$.
+The DUA-D2C method showed a significant improvement with an accuracy of $93.12\%$ over the traditional approach with no subsets, which yielded an accuracy of $92.43\%$.
+The best method in terms of AUC-ROC happens to be the model with even more subsets (5).
 It also seems that the log loss increases a bit when we increase the number of epochs, $E$.
-The effect of dividing the training set into multiple subsets is the most apparent in this dataset compared to all the previous ones.
-AG News, the largest dataset in our study, presents an ideal case for D2C, as each class contains 30,000 samples.
+The effect of using DUA-D2C is the most apparent in this dataset compared to all the previous ones.
+AG News, the largest dataset in our study, presents an ideal case for DUA-D2C, as each class contains 30,000 samples.
 This abundance of data per class allows for substantial division without sacrificing subset size, which maintains the model’s training stability while leveraging the regularizing effect of partitioning and averaging.
-Consequently, every variation of D2C (across all values of $N$ and $E$) demonstrated accuracy gains over the traditional approach, underscoring D2C’s suitability for mitigating overfitting in text classification tasks.
-In summary, D2C excels with large datasets like AG News, where ample samples per class allow the model to benefit from both training on robust subsets and reducing overfitting through partitioning and averaging.
-This validates D2C’s potential for significant impact in Big Data applications, where traditional models often struggle to fully exploit available data.
-VI-D Augmentation in D2C
-Our experiments show that dividing training data into multiple subsets and applying central averaging results in less overfitting and an improvement in performance metrics like accuracy, F1 score, etc.
+Consequently, every variation of DUA-D2C (across all values of $N$ and $E$) demonstrated accuracy gains over the traditional approach, underscoring DUA-D2C’s suitability for mitigating overfitting in text classification tasks.
+In summary, DUA-D2C excels with large datasets like AG News, where ample samples per class allow the model to benefit from both training on robust subsets and reducing overfitting through partitioning and averaging.
+This validates DUA-D2C’s potential for significant impact in Big Data applications, where traditional models often struggle to fully exploit available data.
+Table 8. Comparison of Best Performance: Traditional vs. D2C vs. DUA-D2C across Datasets for Key Metrics.
+Dataset Metric Traditional D2C (Best) DUA-D2C (Best)
+CIFAR-10 Accuracy 0.8604 0.8613 0.8652
+F1-Score 0.8601 0.8611 0.8648
+ROC-AUC 0.98731 0.98704 0.98761
+Log Loss 0.7198 0.5733 0.5876
+Fashion MNIST Accuracy 0.9312 0.9338 0.9344
+F1-Score 0.9313 0.9337 0.9340
+ROC-AUC 0.99546 0.99600 0.99600
+Log Loss 0.3492 0.2341 0.2314
+MNIST Accuracy 0.9952 0.9959 0.9963
+F1-Score 0.9952 0.9959 0.9963
+ROC-AUC 0.99995 0.99999 0.99999
+Log Loss 0.0258 0.0161 0.0148
+FER2013 Accuracy 0.6498 0.6567 0.6537
+F1-Score 0.6420 0.6465 0.6457
+ROC-AUC 0.89934 0.90320 0.90528
+Log Loss 2.0112 1.3941 1.1934
+Audio_Combined Accuracy 0.6953 0.6942 0.7010
+F1-Score 0.6870 0.6909 0.6973
+ROC-AUC 0.96570 0.96553 0.96560
+Log Loss 1.3419 1.0033 1.0571
+AG News Accuracy 0.9243 0.9295 0.9312
+F1-Score 0.9243 0.9296 0.9312
+ROC-AUC 0.98764 0.98916 0.98905
+Log Loss 0.2611 0.2353 0.2378
+  • Values in D2C (Best) and DUA-D2C (Best) columns represent the optimal performance across all tested N (Number of Subsets) and E (Local Epochs) configurations for that method.
+Bold indicates the best overall performance among the three approaches (Traditional, D2C, DUA-D2C) for each metric and dataset.
+For Log Loss, lower values are better.
+The results presented in Table 8 demonstrates the advantages of the ”divide and conquer” strategy.
+The D2C method consistently outperforms the traditional monolithic training approach across various datasets and key metrics, underscoring the inherent benefits of partitioned training and parameter averaging for improved generalization.
+Building upon this foundation, the DUA-D2C method further refines and perfects this technique.
+Through its dynamic, uncertainty-aware aggregation, DUA-D2C consistently achieves superior performance compared to both the traditional method and the standard D2C.
+While the D2C method itself provides substantial gains, the intelligent weighting scheme of DUA-D2C offers an additional performance enhancement.
+This positions DUA-D2C as the more advanced and robust implementation of our proposed framework, consistently delivering strong, and in many cases, the best, generalization capabilities.
+5.4. Augmentation in DUA-D2C
+Our experiments show that applying DUA-D2C results in less overfitting and an improvement in performance metrics like accuracy, F1 score, etc.
 However, we also saw that the training subsets need to be sufficiently large so that each of the subsets can provide enough training diversity to the edge models, otherwise, the performance suffers a drop.
@@
 This approach ensures the separation of data in each of the subsets as well as the expansion of training data, providing enough samples for all the subsets, thus improving the overall performance.
-We tested this augmentation approach on the CIFAR-10 dataset for $3$, $5$ $7$ subsets.
-The augmentation methods we used were rotation, horizontal flip, and shifting along both axes.
+We tested this augmentation approach on the CIFAR-10 dataset for $3$, 55 & $7$ subsets.
+The augmentation methods we used were simple rotation, horizontal flip, and shifting along both axes.
 Let us take a look at the accuracies we achieved in each case and compare them with the accuracies we got before without applying augmentation.
-Figure 15: Test accuracies for different numbers of subsets, N, using the D2C method on the CIFAR-10 dataset with and without Augmentation.
+Figure 14. Test accuracies for different numbers of subsets, N, using the DUA-D2C method on the CIFAR-10 dataset with and without Augmentation.
 As we can see from Figure 14, the accuracy improved significantly in each case after we applied data augmentation to the subsets of the training data.
-The accuracy increased from $86.13\%$ to $88.73\%$ for $3$ subsets.
-In the case of $5$ subsets, the improvement was even more, reaching $88.12\%$ from just $85.52\%$.
-In the case of $7$ subsets too, the accuracy increased from $85.39\%$ to $87.58\%$.
-This shows that data augmentation techniques can be successfully applied when using the D2C method, and they can effectively treat the problems that arise when the size of the subsets becomes too small.
-VII Limitations
-We have shown that our proposed D2C method treats overfitting and shows an improvement over the traditional approach in multiple domains and datasets.
+The accuracy increased from $86.52\%$ to $88.74\%$ for $3$ subsets.
+In the case of $5$ subsets, the accuracy reached $88.12\%$ from just $86.14\%$.
+In the case of $7$ subsets too, the accuracy increased from $85.56\%$ to $87.68\%$.
+This shows that data augmentation techniques can be successfully applied when using the DUA-D2C method, and they can effectively treat the problems that arise when the size of the subsets becomes too small.
+6. Limitations
+We have shown that our proposed DUA-D2C method treats overfitting and shows an improvement over the traditional approach in multiple domains and datasets.
 However, all these come at a cost.
-Dividing the training set into multiple subsets and training them all separately means repeating the whole training process $N$ times ($N$ = number of subsets), albeit with less data.
+Dividing the training set into multiple subsets and training each subset separately means repeating the entire training process $N$ times ($N$ = number of subsets), albeit with less data.
 This results in an increased training time.
 Let us take a look at the following Figure 15 for MNIST.
-Figure 16: Time elapsed per epoch for the traditional approach and different numbers of subsets, N, using the D2C method on the MNIST dataset.
+Figure 15. Time elapsed per epoch for the traditional approach and different numbers of subsets, N, using the DUA-D2C method on the MNIST dataset.
 The training was done using a V100 GPU in each of these cases for fair comparisons.
@@
 For example, the time taken to train for one epoch was $12.5$ seconds in the case of the traditional implementation.
-When the number of subsets, $N$, was $11$, it took $40$ seconds to train for one global epoch.
+When the number of subsets, $N$, was $11$, it took $43$ seconds to train for one global epoch.
 The training time increased significantly, but it was nowhere near $11$ times the time taken in the traditional approach.
@@
 It is important because such an abrupt increase in training time could mean a big issue when it comes to applying the Divide2Conquer method.
-Lastly, it’s important to note that even if the training time is more when using the D2C method, it doesn’t make any difference at all in the post-training deployment stage, as the edge models (like the one used in the traditional approach) and the central model have the same number of weights, architecture, and computational complexity.
-So, clearly, even if the Training time increases while using the Divide2Conquer method, the Inference time remains exactly the same as the traditional approach.
-VIII Conclusion
-Our study demonstrates that dividing training data into multiple subsets and averaging the weights iteratively can treat overfitting and produce significant performance gains through stability and better generalization.
-In this paper, we applied our proposed Divide2Conquer method and tested it on multiple datasets from different domains.
-This method outperformed the traditional approach in all the datasets in terms of key metrics and showed significantly better resistance against overfitting.
-The D2C method also showed a consistent ability to mitigate the increasing trend of validation loss, and to produce smoother decision boundaries.
-These results prove our hypothesis of this method being able to help generalize better and resist overfitting to be true.
-Moreover, the Divide2Conquer method brings even more improvement when the dataset is larger, outlining its promise in Big Data scenarios, especially since D2C method can be applied on top of other generalization techniques.
-Thus, the proposed D2C method opens the door for future research and experiments using this in several domains.
-Acknowledgments
+However, the inherent parallelism of the D2C/DUA-D2C framework lends itself well to distributed computing environments; deploying edge model training across multiple GPUs or compute nodes could substantially mitigate this increase in training time.
+Lastly, it’s important to note that even if the training time is more when using the DUA-D2C method, it doesn’t make any difference at all in the post-training deployment stage, as the edge models (like the one used in the traditional approach) and the central model have the same number of weights, architecture, and computational complexity.
+So, clearly, even if the Training time increases while using the Divide2Conquer technique, the Inference time remains exactly the same as the traditional approach.
+7. Related works
+Several authors have addressed the issue of overfitting while performing various tasks.
+M.
+Cogswell et al.
+proposed a regularizer called DeCov, which helps reduce overfitting in deep neural networks by Decorrelating Representations (cogswell2015reducing, 5).
+Dropout (srivastava2014dropout, 27) was proposed by Srivastava et al.
+back in 2014, and since then, this technique has been extensively used in very complex neural network architectures successfully.
+It was indeed an outstanding contribution, specifically for deep learning-based models.
+Batch Normalization (ioffe2015batch, 10) proposed by Ioffe and Sergei primarily focuses on better convergence and somewhat contributes to reducing overfitting.
+J.
+Kolluri and V.K Kotte came up with L¼ regularization to solve the problems faced by L1 and L2 regularization techniques (kolluri2020reducing, 12).
+Ensembles are also often used to improve generalization.
+The ensemble-based ELVD model (shanmugavel2023novel, 25) managed to outperform the traditional VGGNet and DropoutNet models in terms of reducing overfitting.
+Zehong Zeng et al.
+came up with an ensemble framework that incorporates several techniques to prevent overfitting (zeng2023ensemble, 32).
+Min-Gu Kim et al.
+also proposed parallel ensemble networks to reduce overfitting in ECG data and prevent the degradation of generalization performance as the training progresses (kim2020ensemble, 11).
+We implemented a method based on federated optimization preliminarily for facial expression recognition using FedNet(siddiqui2022fednet, 26).
+This model achieves excellent results in terms of generalization on the CK+ dataset.
+The K-Fold-Cross-Validation method for evaluation proposed by Korjus K.
+et al.
+(korjus2016efficient, 14) has some similarities to our proposed method in the sense that it also creates subsets by partitioning the dataset and makes use of the whole training data.
+However, this method is different from ours since the erroneous samples still make it to the training phase (K-1/K)-th of the time in K-Fold CV.
+The proposed DUA-D2C method partitions data, and the whole training process is partitioned too, which means none of the subsets is fed with a particular outlier while training bar one, minimizing the impact of these erroneous samples.
+Generalization has also been studied in advanced architectures like GCNs.
+Complexity is preserved while introducing measures such as norm regularization and feature projection, as shown by Giovanni Donghi et al.
+(randgcn2024, 6).
+Recent theoretical advances have deepened our understanding of generalization in deep learning.
+Masegosa and Ortega (masegosa2025pacchernoff, 20) proposed PAC-Chernoff bounds to characterize generalization in the interpolation regime, revealing how smoothness arising from regularization, data augmentation, and architectural invariance can lead to more reliable learning.
+We find this framework particularly compelling, as it provides a rigorous lens through which to interpret the benefits of our uncertainty-aware aggregation in DUA-D2C.
+By dynamically favoring models with smoother and more confident predictions, our method echoes the principles outlined in their theoretical analysis.
+Similarly, Lyu et al.
+(lyu2024generalizationgap, 19) investigated the generalization gap in visual reinforcement learning, introducing a representation distance metric to quantify robustness across environments.
+Although their work focuses on reinforcement learning, it underscores the importance of representation stability, a concept that resonates with our Divide2Conquer philosophy.
+By training models on diverse data partitions and aggregating them based on validation performance, DUA-D2C implicitly promotes representational diversity while mitigating overfitting.
+We deeply appreciate these contributions for their clarity and rigor, and we view DUA-D2C as a practical extension of these theoretical insights into the supervised learning domain.
+Overfitting remains a prevalent issue in supervised machine learning despite methods like Early Stopping, Network Reduction, Training Set Expansion, Regularization, and Dropout being effective (ying2019overview, 30).
+Through DUA-D2C, we build an approach that focuses on achieving even better generalization and can be implemented on top of other overfitting-reducing techniques.
+8. Conclusion
+Our study lays the theoretical groundwork and demonstrates experimentally that our proposed Dynamic Uncertainty-Aware Divide2Conquer (DUA-D2C) method effectively treats overfitting and produces significant performance gains through improved stability and better generalization.
+In this paper, we applied our proposed method and tested it on multiple datasets from different domains.
+This advanced method consistently outperformed both the traditional approach and a simpler D2C averaging scheme in terms of key metrics and showed significantly better resistance against overfitting.
+The DUA-D2C method also showed a consistent ability to mitigate the increasing trend of validation loss and to produce smoother decision boundaries.
+These results prove our hypothesis of this refined method being able to help generalize better and resist overfitting to be true.
+Moreover, the DUA-D2C method often brings even more improvement when the dataset is larger, outlining its promise in Big Data scenarios, especially since the DUA-D2C framework can be readily applied on top of other generalization techniques.
+Thus, the proposed DUA-D2C method opens the door for future research and experiments, furthering the exploration of intelligent distributed learning in several domains.
+Acknowledgements.
 This work was supported by the Institute for Advanced Research (IAR), United International University, under grant UIU-IAR-02-2022-SE-06.
 The authors sincerely appreciate the funding and resources provided, which were essential to the successful execution of this research.
-Appendix A Related Proofs and Derivations
-Derivation of $Var(X)=\mathop{{}\mathbb{E}}(X^{2})-(\mathop{{}\mathbb{E}}(X))^{2}$ : Let the random variable $X$ denote the estimate of $x$.
-The variance of estimating this random variable is defined as the average deviation of $X$ from the expected value of our estimate, $\mathop{{}\mathbb{E}}(X)$, where the deviation is squared for symmetry of difference.
+Availability of Data and Codes
+The datasets analyzed during the study are publicly available. The source codes for this study are available on GitHub at https://github.com/Saiful185/DUA-D2C.
+Appendix A Appendix: Related Proofs and Derivations
+Derivation of $Var(X)=\operatorname{\mathbb{E}}(X^{2})-(\operatorname{\mathbb{E}}(X))^{2}$ : Let the random variable $X$ denote the estimate of $x$.
+The variance of estimating this random variable is defined as the average deviation of $X$ from the expected value of our estimate, $\operatorname{\mathbb{E}}(X)$, where the deviation is squared for symmetry of difference.
 So, the variance can be formulated as:
 $Var(X)$ $=\mathop{{}\mathbb{E}}((X-\mathop{{}\mathbb{E}}(X))^{2})$
-$=\mathop{{}\mathbb{E}}(X^{2}+(\mathop{{}\mathbb{E}}(X))^{2}-2X\mathop{{}\mathbb{E}}(X))$
-=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\mathop{{}\mathbb{E}}(X^{2})+(\mathop{{}\mathbb{E}}(X))^{2}-2\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(\mathop{{}\mathbb{E}}(X))$
-$=\mathop{{}\mathbb{E}}(X^{2})-(\mathop{{}\mathbb{E}}(X))^{2}$
-Here, $(a)$ is due to the linear nature of the expectation operator, and $\mathop{{}\mathbb{E}}(\mathop{{}\mathbb{E}}(X))=\mathop{{}\mathbb{E}}(X)$, because $\mathop{{}\mathbb{E}}(X)$ is not a random variable.
-Derivation of $CoV(XY)=\mathop{{}\mathbb{E}}(XY)-\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$ : Let $X$ and $Y$ be estimations of two random variables $x$ and $y$ respectively.
-The covariance of $X$ and $Y$ is defined as the expected value (or mean) of the product of their deviations from their individual expected values[35].
+$=\operatorname{\mathbb{E}}(X^{2}+(\operatorname{\mathbb{E}}(X))^{2}-2X\operatorname{\mathbb{E}}(X))$
+$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\operatorname{\mathbb{E}}(X^{2})+(\operatorname{\mathbb{E}}(X))^{2}-2\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(X))$
+$=\operatorname{\mathbb{E}}(X^{2})-(\operatorname{\mathbb{E}}(X))^{2}$
+Here, $(a)$ is due to the linear nature of the expectation operator, and $\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(X))=\operatorname{\mathbb{E}}(X)$, because $\operatorname{\mathbb{E}}(X)$ is not a random variable.
+Derivation of $CoV(XY)=\operatorname{\mathbb{E}}(XY)-\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y)$ : Let $X$ and $Y$ be estimations of two random variables $x$ and $y$ respectively.
+The covariance of $X$ and $Y$ is defined as the expected value (or mean) of the product of their deviations from their individual expected values(park2018fundamentals, 22).
 Hence,
-$Cov(X,Y)$ $=\mathop{{}\mathbb{E}}((X-\mathop{{}\mathbb{E}}(X)).(Y-\mathop{{}\mathbb{E}}(Y)))$
-$=\mathop{{}\mathbb{E}}(XY-X\mathop{{}\mathbb{E}}(Y)-Y\mathop{{}\mathbb{E}}(X)+\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y))$
-=$\stackrel{{\scriptstyle\text{(b)}}}{{=}}\mathop{{}\mathbb{E}}(XY)-\mathop{{}\mathbb{E}}(X\mathop{{}\mathbb{E}}(Y))$
-$\qquad-\mathop{{}\mathbb{E}}(Y\mathop{{}\mathbb{E}}(X))+\mathop{{}\mathbb{E}}(\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y))$
-=$\stackrel{{\scriptstyle\text{(c)}}}{{=}}\mathop{{}\mathbb{E}}(XY)-\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$
-$\qquad-\mathop{{}\mathbb{E}}(Y)\mathop{{}\mathbb{E}}(X)+\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$
-$=\mathop{{}\mathbb{E}}(XY)-\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$
+$Cov(X,Y)$ $=\operatorname{\mathbb{E}}((X-\operatorname{\mathbb{E}}(X)).(Y-\operatorname{\mathbb{E}}(Y)))$
+$=\operatorname{\mathbb{E}}(XY-X\operatorname{\mathbb{E}}(Y)-Y\operatorname{\mathbb{E}}(X)+\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y))$
+$\stackrel{{\scriptstyle\text{(b)}}}{{=}}\operatorname{\mathbb{E}}(XY)-\operatorname{\mathbb{E}}(X\operatorname{\mathbb{E}}(Y))$
+$\qquad-\operatorname{\mathbb{E}}(Y\operatorname{\mathbb{E}}(X))+\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y))$
+$\stackrel{{\scriptstyle\text{(c)}}}{{=}}\operatorname{\mathbb{E}}(XY)-\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y)$
+$\qquad-\operatorname{\mathbb{E}}(Y)\operatorname{\mathbb{E}}(X)+\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y)$
+$=\operatorname{\mathbb{E}}(XY)-\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y)$
 […]
+$\operatorname{\mathbb{E}}(XY)$ $=\iint xyf(x,y)dxdy$
 […]
-[…]
-[…]
-=$\stackrel{{\scriptstyle\text{(d)}}}{{=}}\int yf(y)\mathop{{}\mathbb{E}}(x)dy$
-[…]
-$=\mathop{{}\mathbb{E}}(X)E(Y)$
+$=\int yf(y)\int xf(x)dxdy$
+$\stackrel{{\scriptstyle\text{(d)}}}{{=}}\int yf(y)\operatorname{\mathbb{E}}(x)dy$
+$=\operatorname{\mathbb{E}}(X)\int yf(y)dy$
+$=\operatorname{\mathbb{E}}(X)E(Y)$
 […]
 Derivation of $Var(\sum_{i=1}^{k}X_{i})=\sum_{i=1}^{k}Var(X_{i})+\sum_{1\leq i\neq j\leq k}Cov(X_{i},X_{j})$:
-Let us consider the random variables $X$ and $Y$ and arbitrary constants $a,b$ Using LABEL:proof1 we can write,
-$Var(aX+bY)$ $=\mathop{{}\mathbb{E}}((aX+bY)^{2})-(\mathop{{}\mathbb{E}}(aX+bY))^{2}$
-$=(a^{2}\mathop{{}\mathbb{E}}(X^{2})+b^{2}\mathop{{}\mathbb{E}}(Y^{2})+2ab\mathop{{}\mathbb{E}}(XY))$
-$\qquad-(a\mathop{{}\mathbb{E}}(x)+b\mathop{{}\mathbb{E}}(Y))^{2}$
-$=(a^{2}\mathop{{}\mathbb{E}}(X^{2})+b^{2}\mathop{{}\mathbb{E}}(Y^{2})+2ab\mathop{{}\mathbb{E}}(XY))$
-$\qquad-a^{2}(\mathop{{}\mathbb{E}}(X))^{2}-b^{2}(\mathop{{}\mathbb{E}}(Y))^{2}$
-$\qquad-2ab\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$
-=$\stackrel{{\scriptstyle\text{(e)}}}{{=}}a^{2}Var(X)+b^{2}Var(Y)+2abCov(X,Y)$
+Let us consider the random variables $X$ and $Y$ and arbitrary constants $a,b$ . Using LABEL:proof1 we can write,
+$Var(aX+bY)$ $=\operatorname{\mathbb{E}}((aX+bY)^{2})-(\operatorname{\mathbb{E}}(aX+bY))^{2}$
+$=(a^{2}\operatorname{\mathbb{E}}(X^{2})+b^{2}\operatorname{\mathbb{E}}(Y^{2})+2ab\operatorname{\mathbb{E}}(XY))$
+$\qquad-(a\operatorname{\mathbb{E}}(x)+b\operatorname{\mathbb{E}}(Y))^{2}$
+$=(a^{2}\operatorname{\mathbb{E}}(X^{2})+b^{2}\operatorname{\mathbb{E}}(Y^{2})+2ab\operatorname{\mathbb{E}}(XY))$
+$\qquad-a^{2}(\operatorname{\mathbb{E}}(X))^{2}-b^{2}(\operatorname{\mathbb{E}}(Y))^{2}$
+$\qquad-2ab\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y)$
+$\stackrel{{\scriptstyle\text{(e)}}}{{=}}a^{2}Var(X)+b^{2}Var(Y)+2abCov(X,Y)$
 Here, $(e)$ follows directly from LABEL:proof1 and LABEL:proof2. So Using LABEL:proof3 and LABEL:proof4 and generalizing for $k$ random variables we can say:
-$Var(\sum_{i=1}^{k}X_{i})$ $=\sum_{i=1}^{k}Var(X_{i})+\sum_{1\leq i\neq j\leq k}Cov(X_{i},X_{j})$ (18)
+(25) $Var(\sum_{i=1}^{k}X_{i})$ $=\sum_{i=1}^{k}Var(X_{i})+\sum_{1\leq i\neq j\leq k}Cov(X_{i},X_{j})$